### v1

# arXiv:2603.01219v1[math.AT] 01 Mar 2026

Abstract.
Using the notion of isotopy modulo $k$, with $k\in\mathbb{N}^{+}$, we introduce a stratification on the set of all minimal $C_{\infty}$-algebra enhancements of a finite-type graded commutative algebra $H^{*}$. We determine obstruction classes defining the extendability of isotopy modulo $k$ to isotopy modulo $(k+1)$ for minimal $C_{\infty}$-algebra enhancements of $H^{*}$ and demonstrate their generalized additivity. As a result, we define a complete set of invariants of the rational homotopy types of closed simply connected manifolds $M$. We prove that if $M$ is a closed (r1)(r-1)-connected manifold of dimension $n\leq l(r-1)+2$ (where $r\geq 2,l\geq 4$), the real and rational homotopy type of $M$ is defined uniquely by the cohomology algebra $H^{*}(M,\mathbb{F})$ and the isotopy modulo (l2)(l-2) of the corresponding minimal unital cyclic $C_{\infty}$-algebra enhancements of $H^{*}(M,\mathbb{F})$ for $\mathbb{F}=\mathbb{R},\mathbb{Q}$, respectively. Combining this with the Hodge homotopy introduced in [11] and developed in [13], we provide a new proof of a theorem by Crowley–Nordström [10]: a (r1)(r-1)-connected closed manifold $M$ of dimension 4r14r-1 with $b_{r}(M)\leq 3$ is intrinsically formal if there exists a $\varphi\in H^{2r-1}(M,\mathbb{R})$ such that the map $H^{r}(M,\mathbb{R})\to H^{3r-1}(M,\mathbb{R}),x\mapsto\varphi\cup x$ is an isomorphism. Furthermore, we provide a new proof and extension of Cavalcanti’s result [5], showing that a (r1)(r-1)-connected closed manifold $M$ of dimension $4r$ with $b_{r}(M)\leq 2$ is intrinsically formal under similar conditions.
Key words and phrases:
Rational homotopy type, minimal $C_{\infty}$-algebra, isotopy modulo $k$, Poincaré Differential Graded Algebra, Hodge homotopy, filtered Differential Graded Lie Algebras (DGLAs), Harrison cohomology, cyclic cohomology, hard Lefschetz property, formality.
2010 Mathematics Subject Classification:
Primary: 55P62; Secondary: 57R19, 13D03, 17B70
1. Introduction
Algebraic models for topological spaces were pioneered by Sullivan, who used differential graded commutative algebras (DGCAs) of differential forms to define the diffeomorphism type of compact manifolds up to finite ambiguity [37, 36]. The Sullivan DGCA model was refined by Halperin-Stasheff [17]. The Halperin-Stasheff filtered model, built on the Tate–Jozefiak resolution of the cohomology algebra $H^{*}({\mathcal{A}}^{*})$ of a DGCA $\mathcal{A}$$^{*}$ over a field $\mathbb{F}$} of characteristic 0, provides a method to construct Sullivan minimal models for simply connected spaces, allowing a systematic approximation of the rational homotopy type [17]. These models, alongside the Quillen differential graded Lie algebras (DGLA) [34] and the Kadeishvili $C_{\infty}$-algebra model [22], have provided robust tools for understanding homotopy types.
In [35] Schlessinger–Stasheff showed that the Halperin-Stasheff model leads to a representation of the set of homotopy type $(X,i:{\mathcal{H}}\cong H^{*}(X))$ as the quotient of a conical rational algebraic variety $V$ modulo a pro-unipotent group action $G$ [35]. Here $V$ is the space of perturbations of the differential in the Tate–Jozefiak resolution of $H^{*}(X)$ and $G$ is pro-unipotent algebraic group. In this formulation “finding a complete set of invariants for rational homotopy types seems to be about the same order of difficulty as finding a complete set of invariants for the $G$-orbits of a variety VV… except for special cohomology rings” [35]. Recent refinements using unital cyclic $C_{\infty}$-algebras by Kajiura [25, 26] and Hamilton–Lazarev [18] have further specialized these tools for closed oriented manifolds. Significant results include Miller’s 1979 formality result for (k1)(k-1)-connected closed manifolds of dimension less than or equal 4k24k-2 [32], using Quillen theory, and Crowley–Nordström’s 2021 classification of (r1)(r-1)-connected manifolds of dimension less than or equal (5r3)(5r-3) via the Bianchi–Massey tensor, using Sullivan theory [10].
This paper aims to refine Kadeishvili RHT classification using minimal $C_{\infty}$-algebras to define a complete set of invariants of the rational homotopy type of a topological space whose cohomology algebra is of finite type. Kadeishvili results [20], [22] stated that for a DGCA $\mathcal{A}$$^{*}$ there exists a minimal $C_{\infty}$-algebra enhancement $C_{\infty}{\mathcal{A}}^{*}$ of $H^{*}({\mathcal{A}}^{*})$ which is homotopy equivalent to $\mathcal{A}$$^{*}$. Furthermore, $C_{\infty}{\mathcal{A}}^{*}$ and $C_{\infty}{\mathcal{B}}^{*}$ are isotopic if and only if $\mathcal{A}$$^{*}$ and $\mathcal{B}$$^{*}$ are weakly equivalent.
We introduce a stratification of isotopy modulo $k$, for all $k\in{\mathbb{N}}^{+}$, on the set of all minimal $C_{\infty}$-algebra enhancements of a finite type graded commutative algebra (GCA) $H^{*}$ and show that two such $C_{\infty}$-algebra enhancements are isotopic if and only if they are isotopic modulo $k$ for any $k\in{\mathbb{N}}^{+}$. As a result, we give classifications of real and rational homotopy types of connected, simply connected smooth manifolds $M$ with the same cohomology algebra $H^{*}(M,{\mathbb{F}})$, $\mathbb{F}$$\mathbb{Q}$$\mathbb{R}$}, in terms of a sequence of obstruction classes defining the extendability of isotopy modulo $k$ to isotopy modulo $(k+1)$ of minimal $C_{\infty}$-algebra enhancements of $H^{*}(M,{\mathbb{F}})$. We show that the real and rational homotopy types of a closed (r1)(r-1)-connected manifold of dimension $n\leq l(r-1)+2$, $r\geq 2$, $l\geq 4$, is defined uniquely by the isotopy modulo (l2)(l-2) of the corresponding minimal unital cyclic $C_{\infty}$-algebra enhancements of $H^{*}(M,{\mathbb{F}})$, $F={\mathbb{R}}$, $\mathbb{Q}$}, respectively (Theorem 4.17). Furthermore, we illustrate the technique of Hodge homotopy, introduced by Fiorenza–Kawai–Lê–Schwachhöfer in [11], and developed further by Fiorenza–Lê in [13], to give a new proof of a formality theorem for (r1)(r-1)-connected closed manifolds $M$ of dimension 4r14r-1 due to Crowley–Nordström [10] and an extension of a formality theorem for (r1)(r-1)-connected closed manifolds $M$ of dimension $4r$ due to Cavalcanti [5].
Our paper is organized as follows. In Section 2, we recall the concepts of a Hodge homotopy $d^{-}$, the corresponding harmonic subspace, and the Hodge decomposition, which is stronger than those considered by Chuang-Lazarev [7] and by Kajiura [25]. In Section 3, we show that the transferred minimal $C_{\infty}$-algebra associated with a Hodge homotopy $d^{-}$ is unital cyclic (Theorem 3.3). We also consider several symmetries of the minimal unital cyclic $C_{\infty}$-algebra, transferred using Hodge homotopy, that shall be needed in the proof of formality theorems in later sections. Section 4 is devoted to classification of real-and rational homotopy type of simply connected topological spaces, in particular, closed simply connected manifolds, using minimal $C_{\infty}$-algebra and minimal unital cyclic $C_{\infty}$-algebra enhancements and their isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, deformation theory of filtered differential graded Lie algebras (DGLAs), Harrison cohomology, and cyclic cohomology. We give explicit formulas for the first and secondary obstruction classes (Proposition 4.8, Lemmas 4.9, 4.10, Proposition 4.11). We describe the generalized affine structure of the set of complete invariants of the rational homotopy types of simply-connected topological spaces with a given cohomological algebra of finite type (Theorem 4.12, Proposition 4.16). Then we give a proof of Theorem 4.17 on the rational homotopy type of a (r1)(r-1)-connected, $r\geq 2$, manifold of dimension less than or equal l(r1)+2l(r-1)+2, $l\geq 4$. We illustrate our technique in Subsection 4.5. In Sections 5 and 6, we give proofs of formality and almost formality theorems (Theorems 5.1, 6.1), using Hodge homotopy and the resulting symmetries of the transferred minimal unital cyclic $C_{\infty}$-algebras, proved in Section 3. In the last section we discuss related results and some open questions.
The results presented in this paper concerning the case of (r1)(r-1)-connected manifolds of dimension n=4r1n=4r-1 with $b_{r}\leq 3$, as well as several foundational technical lemmas in the preliminaries, were developed in collaboration with Domenico Fiorenza. The author is deeply grateful for his remark about a parallel between Hamilton–Lazarev’s obstruction theory and the theory we develop in this paper, and for his generous guidance. The proofs regarding the $n=4r$ case, the higher-dimensional bounds up to 5r25r-2 and l(r1)+2l(r-1)+2, the introduction of isotopy modulo $k$ and the resulting stratification on the set of minimal $C_{\infty}$-algebra enhancements, the gauge-theoretic analysis of the higher obstruction classes are the work of the current author.
2. Hodge homotopy and unital cyclic $C_{\infty}$-algebras
In this section we briefly recall the concept of a Poincaré DGCA (PDGCA) admitting a Hodge homotopy, which was introduced by Fiorenza–Kawai–Lê-Schwachhöfer in [11], and developed further by Fiorenza–Lê in [13]. We show that Poincaré DGCAs are unital cyclic $C_{\infty}$-algebras (Remark 2.6. (3+4)).
Given a vector space $V$ over a field $\mathbb{F}$} we denote by $V^{\vee}$ the dual space of $V$.
Definition 2.1.
[13, Def. 2.1], cf. [28, Def. 4.1], cf. [11, Definition 2.1]. A PDGCA $({\mathcal{A}}^{*},d)$ of degree $n$ over a field $\mathbb{F}$} is a DGCA ${\mathcal{A}}^{*}=\oplus_{k=0}^{n}{\mathcal{A}}^{k}$ whose cohomology ring $H^{*}({\mathcal{A}})$ is finite dimensional over $\mathbb{F}$} and has a fundamental class $\int\in(H^{n}({\mathcal{A}}))^{\vee}$ such that the pairing, defined by using the cup-product,
(2.1) $\langle\alpha^{k},\beta^{l}\rangle=\begin{cases}\int\alpha^{k}\cdot\beta^{l}&\text{ if }k+l=n,\\ 0&\text{ else, }\end{cases}$
is non-degenerate, i.e. $\langle\alpha,H^{*}({\mathcal{A}})\rangle=0$ if and only if $\alpha=0$.
The pairing on $H^{*}({\mathcal{A}})$ induces a pairing on $\mathcal{A}$$^{*}$ as follows
(2.2) $\langle\alpha^{k},\beta^{l}\rangle=\begin{cases}\int[\alpha^{k}\cdot\beta^{l}]&\text{ if }k+l=n,\\ 0&\text{ else, }\end{cases}$
where $\cdot$] stands for the projection ${\mathcal{A}}^{n}\subseteq\ker d\to H^{n}({\mathcal{A}}^{*})$. The PDGCA $({\mathcal{A}}^{*},d)$ is called non-degenerate, [11, Definition 2.2], or oriented Poincaré duality algebra [28, Definition 2.1] if the pairing $\langle\cdot,\cdot\rangle$ on $\mathcal{A}$$^{*}$ is non-degenerate. A PDGCA $(H^{*},d=0)$ is called a Poincaré GCA.
Definition 2.2.
[13, Definition 2.2] Let $({\mathcal{A}}^{\ast},d,\cdot,\langle-,-\rangle)$ be a Poincaré DGCA. A Hodge homotopy for $({\mathcal{A}}^{\ast},d,\cdot,\langle-,-\rangle)$ is a degree 1-1 operator $d^{-}\colon{\mathcal{A}}^{\ast}\to{\mathcal{A}}^{\ast-1}$ such that the following orthogonality relations hold:
(2.3) $d^{-}d^{-}=0;\quad d^{-}dd^{-}=d^{-};\quad dd^{-}d=d;$
(2.4) $\langle\mathrm{Im}(d^{-}),\mathrm{Im}(d^{-})\rangle=0;\quad\langle\mathrm{Im}(\pi_{\mathcal{H}^{\ast}}),\mathrm{Im}(d^{-})\rangle=0,$
where
(2.5) $\pi_{\mathcal{H}^{\ast}}:=\mathrm{id}_{{\mathcal{A}}^{\ast}}-[d,d^{-}].$
Remark 2.3.
[13, Remark 2.3] It follows from (2.3) that the two endomorphisms $dd^{-},d^{-}d\colon{\mathcal{A}}^{\ast}\to{\mathcal{A}}^{\ast}$ are orthogonal idempotents. So they both, and also $\pi_{\mathcal{H}}$, are projection operators.
Definition 2.4.
[11, Definition 2.2], [13, Definition 2.4]. The graded subspace $\mathcal{H}^{\ast}:=\mathrm{Im}(\pi_{\mathcal{H}^{\ast}})$ is called the harmonic subspace for the Hodge homotopy $d^{-}$.
Remark 2.5.
[13, Remark 2.5] It follows from the definition that $d\pi_{\mathcal{H}^{\ast}}=\pi_{\mathcal{H}^{\ast}}d=d^{-}\pi_{\mathcal{H}^{\ast}}=\pi_{\mathcal{H}^{\ast}}d^{-}=0$, and that we have a direct sum decomposition
(2.6) ${\mathcal{A}}^{\ast}=\mathcal{H}^{\ast}\oplus\underbrace{dd^{-}{\mathcal{A}}^{\ast}\oplus d^{-}d{\mathcal{A}}^{\ast}}_{{\mathcal{L}_{\mathcal{A}}^{\ast}}}$
with $({\mathcal{L}_{\mathcal{A}}^{\ast}},d)$ an acyclic complex and $d^{-}({\mathcal{L}_{\mathcal{A}}^{\ast}})\subseteq{\mathcal{L}_{\mathcal{A}}^{\ast}}$. Also, the composition
$\mathcal{H}^{\ast}\to\ker(d)\to H^{\ast}({\mathcal{A}}^{*})$
is an isomorphism of graded vector spaces.
Remark 2.6.
(1) It follows from (2.2) that the Poincaré pairing is graded symmetric, i.e., for all $x,y\in{\mathcal{A}}^{*}$ we have
(2.7) $\langle x,y\rangle=(-1)^{|x||y|}\langle y,x\rangle.$
(2) Since $\int d\omega=0$, $d$ is graded skew-adjoint, i.e., for all $x,y\in{\mathcal{A}}^{*}$ we have
(2.8) $\langle dx,y\rangle=-(-1)^{|x|}\langle x,dy\rangle.$
This implies that the direct sum decomposition ${\mathcal{A}}^{\ast}=\mathcal{H}^{\ast}\oplus{\mathcal{L}_{\mathcal{A}}^{\ast}}$ is a orthogonal direct sum decomposition: ${\mathcal{A}}^{\ast}=\mathcal{H}^{\ast}\oplus^{\perp}{\mathcal{L}_{\mathcal{A}}^{\ast}}$.
(3) It follows from (2.2) that $\langle-,$-$\rangle$ is an invariant inner product:
(2.9) $\langle xy,z\rangle=\langle x,yz\rangle\>\>\forall x,y,z\in{\mathcal{A}}^{*}.$
Writing $dx=\mu_{1}(x)$ and $xy=\mu_{2}(x,y)$ equations (2.8) and (2.9) can be written as
$\langle\mu_{1}(x),y\rangle=(-1)^{1}(-1)^{(|x||y|)}\langle\mu_{1}(y),x\rangle$
and
$\langle\mu_{2}(x,y),z\rangle=(-1)^{|x|(|y|+|z|)}\langle\mu_{2}(y,z),x\rangle,$
i.e., the DGCA $(\mathcal{A}^{*},d)$ seen as a $C_{\infty}$-algebra with $\mu_{k}=0$ for $k>2$ is a cyclic $C_{\infty}$-algebra in the sense of [7] with the symmetric group action on operations given by $\mu_{k}^{\sigma}=(-1)^{|\sigma|}\mu_{k}$, where $|\sigma|$ is the sign of the permutation. In particular, for the cyclic permutation $z(k)=(0,1,\dots,k)\in\mathbb{S}_{k+1}$ acting on $\mu_{k}$ we find $\mu_{k}^{z(k)}=(-1)^{k}\mu_{k}$.
(4) We recall that a unital $C^{\infty}$-algebra is a $C^{\infty}$-algebra with a unit $1$ of degree 0 such that $d(1)=0$, $m_{2}(1,a)=m_{2}(a,1)$, and for $k\geq 3$ we have $m_{k}(a_{1},\ldots,1,\ldots,a_{k})=0$ for all $a,a_{1},\ldots a_{k}\in{\mathcal{A}}^{*}$ [6, Definition 8]. Clearly a Poincaré GCA $H^{*}$ is unital and the harmonic representative of the unit $1\in H^{0}$ is the unit $1\in\mathcal{A}^{0}$.
Examples 2.7.
1) Our working example of a PDGCA over $\mathbb{R}$} admitting a Hodge homotopy is the de Rham complex $({\Omega}^{*}(M),d)$ of a closed smooth Riemann manifold $M$ with the harmonic subspace being the space of harmonic forms on $M$ [11, Remark 2.7 (1)], [13, Example 2.7]. In [11, Theorem 3.10, (4.1)] Fiorenza-Kawai–Lê–Schwachhöfer proposed a simple algorithm to find a nondegenerate finite dimensional PDGCA which is weakly equivalent to $({\Omega}^{*}(M),d)$ if $b_{1}(M)=0$.
2) Our working example of a PDGCA over $\mathbb{Q}$} admitting a Hodge homotopy is the finite dimensional Poincaré duality algebra which is weakly equivalent to a given simply connected PDGCA over $\mathbb{Q}$} by [28, Theorem 1.1]. By [8, Lemma 11.1] such a finite dimensional Poincaré duality algebra admits a Hodge homotopy.
3. Cyclicity of the minimal unital $C_{\infty}$-algebra via Hodge homotopy transfer
Let $(A^{*},d,\cdot,\langle-,-\rangle)$ be a PDGCA of degree $n$ over a field $\mathbb{F}$}, equipped with a Hodge homotopy $d^{-}$. Note that we don’t require that $\langle-,$-$\rangle$ is nondegenerate. Moreover, we don’t assume that $H^{1}({\mathcal{A}}^{*})=0$. We show that the unital minimal $C_{\infty}$-enhancement of $\big({\mathcal{H}}^{*},d=0\big)$ obtained from $(\mathcal{A}^{*},d,\cdot,\langle-,-\rangle)$ by homotopy transfer via $d^{-}$ is cyclic (Theorem 3.3).
Lemma 3.1.
We have
  1. (1)
    $j$ is an isometric embedding,
  2. (2)
    For all $x\in{\mathcal{H}},y\in{\mathcal{A}}^{*}$ we have
    (3.1) $\langle x,y\rangle=\langle x,\pi_{\mathcal{H}}y\rangle.$
  3. (3)
    $d^{-}$ is graded self-adjoint:
    (3.2) $\langle d^{-}x,y\rangle=(-1)^{|x|}\langle x,d^{-}y\rangle.$
Proof.
The first and second assertion of Lemma 3.1 are consequences of the Hodge decomposition (2.6). To prove the last assertion, we use the identity $\mathrm{id}=\pi_{\mathcal{H}}+dd^{-}+d^{-}d$ together with graded skew-adjointness of $d$ (2.8), and the orthogonality relations:
$\langle d^{-}x,y\rangle=\langle d^{-}x,dd^{-}y\rangle=(-1)^{|x|}\langle dd^{-}x,d^{-}y\rangle=(-1)^{|x|}\langle x,d^{-}y\rangle$
This completes the proof of Lemma 3.1. ∎
Remark 3.2.
By (2.5), we have that
${\mathcal{H}}^{*}\xrightarrow{j}{\mathcal{A}}^{*}\xrightarrow{\pi_{{\mathcal{H}}^{*}}}{\mathcal{H}}^{*},\qquad d^{-}:{\mathcal{A}}^{*}\to{\mathcal{A}}^{*-1},$
is a contraction datum for $(\xrightarrow{j}{\mathcal{A}}^{*},d)$. Namely, we have
$\pi_{\mathcal{H}}\circ j=\mathrm{Id},\qquad j\pi_{\mathcal{H}}-id=dd^{-}+d^{-}d.$
We can therefore use homotopy transfer to transfer the unital cyclic DGCA structure of $({\mathcal{A}}^{*},d)$ to a $C_{\infty}$-structure on $\mathcal{H}$$^{*}$. This structure will be a minimal one since the differential on $\mathcal{H}$$^{*}$ is zero.
Theorem 3.3.
Then the transferred minimal $C_{\infty}$-algebra associated with a Hodge homotopy $d^{-}$, denoted by $C_{\infty}({\mathcal{H}}^{*})$, is unital and cyclic with the symmetric group action on operations given by the sign representation, i.e., for $x_{1},\dots,x_{k+1}\in\mathcal{H}^{*}$ we have
(3.3) $\langle m_{k}(x_{1},\dots,x_{k}),x_{k+1}\rangle=(-1)^{k}(-1)^{\epsilon}\langle m_{k}(x_{2},\dots,x_{k+1}),x_{1}\rangle,$
where $\epsilon=|x_{1}|(|x_{2}|+\dots+|x_{k+1}|)$.
Proof of Theorem 3.3.
Since the Poincaré pairing is graded symmetric (2.7), by Lemma 3.1, the Poincaré pairing $\langle-,$-$\rangle$, the projection $\pi_{\mathcal{H}}$, and the homotopy $d^{-}$ satisfy the conditions for a harmonic Hodge decomposition in [7, Definition 2.1]. Hence, by [7, Theorem 4.2] the transferred minimal $C_{\infty}$-algebra is cyclic. Moreover, since the DGCA structure on $\mathcal{A}$$^{*}$ is a particular instance of a unital $C_{\infty}$-algebra with unit $1\in\mathbb{K}={\mathcal{A}}^{0}$, and the homotopy $d^{-}$ satisfies side conditions [6, Theorems 10, 12], i.e.,
(3.4) $\pi_{{\mathcal{H}}^{*}}\circ{d^{-}}=0,\;({d^{-}})^{2}=0,\>{d^{-}}(1)=0,$
we have that the transferred cyclic minimal $C_{\infty}$-algebra $C_{\infty}({\mathcal{H}}^{*})$ is unital [6, Theorems 10, 12]. ∎
Remark 3.4.
Theorem 3.3 can be also obtained from [25, Lemma 5.23], where Kajiura used a Hodge-Kodaira decomposition to obtain a minimal cyclic $A_{\infty}$-algebra from a cyclic $A_{\infty}$-algebra. Note that our Hodge decomposition (2.6) is also a particular case of the Hodge-Kodaira decomposition defined in [25, Section 5.2].
Remark 3.5.
Let $r\geq 1$ and let $({\mathcal{A}}^{*},d,\langle,-,-\rangle)$ be a (r1)(r-1)-connected Poincaré DGCA of degree $n$ admitting a Hodge homotopy $d^{-}$ whose Hodge subspace is $\mathcal{H}$$^{*}$. The explicit expression for $m_{k}$ of $C_{\infty}({\mathcal{H}}^{*})$ are derived by using [31, Theorem 3.4], [6, Theorem 12]. One considers the recursive formula
$\widehat{m}_{2}(\alpha_{1},\alpha_{2})$ $=\alpha_{1}\cdot\alpha_{2},$
$\widehat{m}_{k}(\alpha_{1},$ $\dots,\alpha_{k})=(-1)^{k-1}{d^{-}}\widehat{m}_{k-1}(\alpha_{1},\dots,\alpha_{k-1})\cdot\alpha_{k}$
$-(-1)^{k\deg(\alpha_{1})}\alpha_{1}\cdot{d^{-}}\widehat{m}_{k-1}(\alpha_{2},\cdots,\alpha_{k})$
(3.5) $-\sum_{i=2}^{k-2}(-1)^{\nu}{d^{-}}\widehat{m}_{i}(\alpha_{1},\dots,\alpha_{i})\cdot{d^{-}}\widehat{m}_{k-i}(\alpha_{i+1},\dots,\alpha_{k}),$
where the $\alpha_{i}$’s are element in ${\mathcal{H}}^{\ast}\subseteq{\mathcal{A}}^{\ast}$ and $\cdot$ is the multiplication in the DGCA $({\mathcal{A}}^{\ast},d)$, and where $\nu=i+(k-i-1)(\deg(\alpha_{1})+\cdots+\deg(\alpha_{i}))$. Then the multiplication $m_{k}$ is defined by
(3.6) ${m}_{k}(\alpha_{1},\dots,\alpha_{k})=\pi_{{\mathcal{H}}^{*}}\left(\widehat{m}_{k}(\alpha_{1},\dots,\alpha_{k})\right).$
For instance, the transferred ternary multiplication $m_{3}$ is explicitly given by:
$m_{3}(x,y,z)$ $=\pi_{{\mathcal{H}}^{*}}\big(d^{-}(x\cdot y)\cdot z\big)$
(3.7) $\qquad-(-1)^{|x|}\pi_{{\mathcal{H}}^{*}}\big(x\cdot d^{-}(y\cdot z)\big).$
One easily sees inductively that the $m_{k}$ are unital, i.e., that $m_{k}(\alpha_{1},\dots,\alpha_{k})$ vanishes if one of the $\alpha_{i}$’s is equal to $1$. In [13, Remark 2.15] Fiorenza-Lê showed that for any $k\geq 3$ and for any $\alpha\in\mathcal{H}^{r}$, one has the following vanishing property:
(3.8) $\widehat{m}_{k}(\alpha,\alpha,\dots,\alpha)=m_{k}(\alpha,\alpha,\dots,\alpha)=0.$
Lemma 3.6 (Generalized Odd Symmetry).
Assume that $\mathcal{A}$$^{*}$ is (r1)(r-1)-connected, with $r$ odd. Then for every $x,y,z\in{\mathcal{H}}^{r}$ we have
(3.9) $\widehat{m}_{k}(x,\ldots,x,y)=(-1)^{k-1}\widehat{m}_{k}(y,x,\ldots,x)$
(3.10) $\widehat{m}_{k}(x,y,\ldots y,z)=(-1)^{k-1}\widehat{m}_{k}(z,y\ldots,y,x)$
Proof.
We prove (3.9) by induction on $k$. For $k=3$, equation (3.9) follows directly from (3.7). Assume that (3.9) holds for all indices up to (k1)(k-1). By (3.5) and by the the vanishing property (3.8) we have
$\widehat{m}_{k}(x,\ldots,x,y)$ $=(-1)^{k-1}d^{-}\widehat{m}_{k-1}(x,\ldots,x)\cdot y$
$\qquad-(-1)^{k}x\cdot d^{-}\widehat{m}_{k-1}(x,\ldots,x,y)$
$\qquad-\sum_{i=2}^{k-2}(-1)^{i(k-i)}d^{-}\widehat{m}_{i}(x,\ldots,x)\cdot d^{-}\widehat{m}_{k-i}(x,\ldots,x,y)$
$=(-1)^{k-1}x\cdot d^{-}\widehat{m}_{k-1}(x,\ldots,x,y)$
$\qquad-d^{-}\widehat{m}_{2}(x,x)\cdot d^{-}\widehat{m}_{k-2}(x,\ldots,x,y).$
and, taking also in account the inductive assumption, and recalling that $r$ is odd,
$\widehat{m}_{k}(y,x,\ldots,x)$ $=(-1)^{k-1}d^{-}\widehat{m}_{k-1}(y,\ldots,x)\cdot x$
$\qquad-(-1)^{k}y\cdot d^{-}\widehat{m}_{k-1}(x,\ldots,x)$
$\qquad-\sum_{i=2}^{k-2}(-1)^{i(k-i)}d^{-}\widehat{m}_{i}(y,x,\ldots,x)\cdot d^{-}\widehat{m}_{k-i}(x,\ldots,x)$
$=(-1)^{k-1}d^{-}\widehat{m}_{k-1}(y,x\ldots,x)\cdot x$
$\qquad-d^{-}\widehat{m}_{k-2}(y,\ldots,x)\cdot d^{-}\widehat{m}_{2}(x,x)$
$=-d^{-}\widehat{m}_{k-1}(x,\ldots,x,y)\cdot x$
$\qquad+(-1)^{k-3}d^{-}\widehat{m}_{k-2}(x,\ldots,x,y)\cdot d^{-}\widehat{m}_{2}(x,x)$
$=x\cdot d^{-}\widehat{m}_{k-1}(x,\ldots,x,y)$
$\qquad+(-1)^{k-2}d^{-}\widehat{m}_{2}(x,x)\cdot d^{-}\widehat{m}_{k-2}(x,\ldots,x,y).$
Equation (3.9) follows. The proof of (3.10) is analogous. ∎
Lemma 3.7 (Generalized Even Symmetry for $k=3$).
Assume that $\mathcal{A}$$^{*}$ is (r1)(r-1)-connected, with $r$ even. Then for every $y\in{\mathcal{H}}^{r}$, for every homogeneous $x,z$ we have
(3.11) $\widehat{m}_{3}(x,x,y)=-(-1)^{|x|}\widehat{m}_{3}(y,x,x)$
(3.12) $\widehat{m}_{3}(x,y,z)=-(-1)^{|x||z|}\widehat{m}_{3}(z,y,x)$
In particular, for every $x,y\in\mathcal{H}^{r}$, we have
(3.13) $\widehat{m}_{3}(x,y,x)=0.$
Proof.
Immediate from equation (3.5). ∎
Lemma 3.8 (Generalized Even Symmetry for $k=4$).
Assume that $\mathcal{A}$$^{*}$ is (r1)(r-1)-connected, with $r$ even. Then for every $x,y,z\in{\mathcal{H}}^{r}$ we have
(3.14) $\widehat{m}_{4}(x,x,x,y)=-\widehat{m}_{4}(y,x,x,x)$
(3.15) $\widehat{m}_{4}(x,y,y,z)=-\widehat{m}_{4}(z,y,y,x)$
Proof.
By (3.5) and by the vanishing condition (3.8), we have
$\widehat{m}_{4}(x,x,x,y)$ $=-d^{-}\widehat{m}_{3}(x,x,x)\cdot y-x\cdot d^{-}\widehat{m}_{3}(x,x,y)$
$\qquad\qquad-d^{-}\widehat{m}_{2}(x,x)\cdot d^{-}\widehat{m}_{2}(x,y)$
$=-x\cdot d^{-}\widehat{m}_{3}(x,x,y)-d^{-}\widehat{m}_{2}(x,x)\cdot d^{-}\widehat{m}_{2}(x,y).$
Similarly, and taking into account (3.11),
$\widehat{m}_{4}(y,x,x,x)$ $=-d^{-}\widehat{m}_{3}(y,x,x)\cdot x-y\cdot d^{-}\widehat{m}_{3}(x,x,x)$
$\qquad\qquad-d^{-}\widehat{m}_{2}(y,x)\cdot d^{-}\widehat{m}_{2}(x,x)$
$=-d^{-}\widehat{m}_{3}(y,x,x)\cdot x-d^{-}\widehat{m}_{2}(y,x)\cdot d^{-}\widehat{m}_{2}(x,x)$
$=d^{-}\widehat{m}_{3}(x,x,y)\cdot x+d^{-}\widehat{m}_{2}(x,x)\cdot d^{-}\widehat{m}_{2}(y,x).$
This proves (3.14). The proof of (3.15) is analogous. ∎
Remark 3.9.
Via the canonical graded vector spaces isomorphisms ${\mathcal{H}}^{\ast}\xrightarrow{\sim}H^{\ast}$ between the space of harmonic forms and the cohomology of $\mathcal{A}$$^{\ast}$, the unital cyclic minimal $C_{\infty}$-algebra structure on $\mathcal{H}$$^{\ast}$ is equivalently a unital cyclic minimal $C_{\infty}$-algebra structure on $H^{\ast}$. By abuse of notation we will denote by the same symbols $m_{k}$ the multiplication on $H^{\ast}$ for this structure. Notice that $m_{2}$ is precisely the multiplication $\cdot$ in cohomology.
4. Classification of rational and real homotopy types of closed simply-connected smooth manifolds
In this section, for $k\in{\mathbb{N}}^{+}$, we introduce the concept of equality modulo $k$ (Definition 4.3) and isotopy modulo $k$ (Definition 4.4, Remark 4.6) for minimal $C_{\infty}$-algebra enhancements of a finite type GCA $H^{*}$ over a field $\mathbb{F}$} of characteristic 0. We show that, for any $k\in{\mathbb{N}}^{+}$, isotopy modulo $k$ defines an equivalence relation on the set $C_{\infty}(H^{*})$ of all minimal $C_{\infty}$-algebra enhancements $(H,m_{1}=0,m_{2},m_{3},\ldots)$ of $H^{*}$ whose multiplication is $m_{2}$. We also denote such a minimal $C_{\infty}$-algebra enhancements as $m=(m_{3},m_{4},\ldots)$, (Proposition 4.5), and two minimal $C_{\infty}$-algebra enhancements are isotopic iff they are isotopic modulo $k$ for all $k\in{\mathbb{N}}^{+}$ (Theorem 4.7). We show that if $H^{*}$ is a PDGCA, the primary obstruction for isotopy modulo 3 of a minimal unital cyclic $C_{\infty}$-algebra enhancement $m=(m_{3},m_{4},\ldots)$ of $H^{*}$ is the cyclic cohomology class $Tm_{3}]\in HC^{4,-1}(H^{*})$ (Proposition 4.8). Then we describe the structure of a complete set of invariants of the set $C_{\infty}(H^{*})$ modulo isotopy (Theorem 4.12, Proposition 4.16). We prove that the real and rational homotopy types of a closed (r1)(r-1)-connected manifolds of dimension $n\leq l(r-1)+2$ (where $r\geq 2,l\geq 4)$ is defined uniquely by the cohmology algebra $H^{*}(M,{\mathbb{Q}})$ and the isotopy modulo l2l-2 of the corresponding $C_{\infty}$-algebra enhancement of $H^{*}(M,{\mathbb{Q}})$ (Theorem 4.17.) Finally we demonstrate our technique on a non-formal 8-dimensional nilmanifold in Subsection 4.5.
We begin with Subsection 4.1 where we fix notation and recall preliminary results.
4.1. $C_{\infty}$-isotopies, twisting cochains, filtered differential graded Lie algebras, and cyclic cohomology
(i) Let us recall that a $C_{\infty}$-morphism $\phi=(\phi_{1},\phi_{2},\phi_{3},\ldots)$ from a minimal $C_{\infty}$-algebra $A=(H^{*},m_{1}=0,m_{2},\ldots)$ to a minimal $C_{\infty}$-algebra $A^{\prime}=(H^{*},m_{1}^{\prime}=0,m_{2}^{\prime},\ldots)$ must satisfy [30, Proposition 10.2.6, p. 374]:
$\sum_{k=1}^{n}\sum_{r_{1}+\ldots r_{k}=n}(-1)^{\eta}m_{k}^{\prime}(\phi_{r_{1}}(a_{1},\ldots,a_{r_{1}}),\ldots,\phi_{r_{k}}(a_{n-r_{k}+1},\ldots,a_{n}))$
(4.1) $=\sum_{k=1}^{n}\sum_{\lambda=0}^{n-k}(-1)^{\xi}\phi_{n-k+1}(a_{1},\ldots,a_{\lambda},m_{k}(a_{\lambda+1},\ldots,a_{\lambda+k}),a_{\lambda+k+1},\ldots,a_{n})$
for all $n\geq 2$ and
$\eta=\sum_{1\leq\alpha<\beta\leq k}(|a_{{r_{1}}+\ldots+r_{\alpha-1}+1}|+\ldots+|a_{{r_{1}}+\ldots+_{r_{\alpha}}}|+r_{\alpha})(1+r_{\beta}),$
$\xi=k+k\lambda+k(|a_{1}|+\ldots+|a_{\lambda}|)+n.$
We call $A$ and $A^{\prime}$ isotopic if $\phi=(\phi_{1}=\mathrm{Id},\phi_{2},\phi_{3},\ldots)$, i.e., $\phi$ is a $C_{\infty}$-isotopy.
(ii) Our notion of isotopy modulo $k$ on the set $C_{\infty}(H^{*})$ in Definition 4.4 is motivated by Kadeishvili’s theory of twisting cochains for expressing the isotopy defined by the action of $C_{\infty}$-isotopies on the set of minimal $C_{\infty}$-enhancements of a GCA $H^{*}$ of finite type [21], [22, Section 6 6]. Denote by $C^{*,*}(H^{*},H^{*})$ the bigraded Harrison cochain complex on $H^{*}$ with values in $H^{*}$. First we consider the structure of a homotopy $G$-algebra on $C^{*,*}(H^{*},H^{*})$ [21], [16]. Two minimal $C_{\infty}$-enhancements, represented by twisting cochains $m=\sum_{i\geq 3}m_{i}$ and $m^{\prime}=\sum_{i\geq 3}m^{\prime}_{i}$, where $m_{i},m_{i}^{\prime}\in C^{i,2-i}(H^{*},H^{*})$, are equivalent if there exists a perturbation cochain $p=\sum_{i\geq 2}p_{i}\in C^{i,1-i}(H^{*},H^{*})$ such that:
(4.2) $m-m^{\prime}=\delta p+p\cdot p+p\circ m+m^{\prime}\circ p+m^{\prime}\{p,p\}+\ldots$
Let $\delta$ be the Harrison differential $d_{Harr}$ up to sign [38], defined as follows [39, (1.3.4)]:
$(-1)^{n+1}\delta f(a_{1}\otimes\ldots\otimes a_{n})$ $\coloneqq a_{1}f(a_{2}\otimes\ldots\otimes a_{n})$
$+\sum_{i=1}^{n-1}(-1)^{i}f(a_{1}\otimes\ldots\otimes a_{i}a_{i+1}\otimes\ldots\otimes a_{n})$
(4.3) $+(-1)^{n}f(a_{1}\otimes a_{2}\ldots\otimes a_{n-1})a_{n}.$
The dot product $\cdot$, the Gerstenhabers product $\circ$, the brace operation $\{\ldots\}$ are defined on Harrison cochain complex $C^{*,*}(H^{*},H^{*})$ as follows [38], [22, §6], [39] :
$\cdot:C^{n,m}(H^{*},H^{*})\otimes C^{n^{\prime},m^{\prime}}(H^{*},H^{*})\mapsto C^{n+n^{\prime},m+m^{\prime}}(H^{*},H^{*})$
$f\cdot g(a_{1}\otimes\ldots\otimes a_{n+m})\coloneqq(-1)^{nn^{\prime}}f(a_{1}\otimes\ldots\otimes a_{n})\cdot g(a_{n+1}\otimes\ldots a_{n+n^{\prime}}),$
$-\{\ldots\}:C^{n,m}(H^{*},H^{*})\otimes_{t=1}^{i}C^{n_{t},m_{t}}(H,H)\to C^{n+\sum_{t}n_{t}-i,m+\sum_{t}m_{t}}(H^{*},H^{*}),$
$(f;g_{1},\ldots,g_{i})\mapsto f\{g_{1},\ldots,g_{i}\},$
$f\{g_{1},\ldots,g_{i}\}(a_{1}\otimes\ldots\otimes a_{n+n_{1}+\ldots+n_{i}-i})$
$\coloneqq\sum_{k_{1}\leq k_{2}\ldots\leq k_{i}}(-1)^{\varepsilon}f(a_{1}\otimes\ldots\otimes a_{k_{1}}\otimes g_{1}(a_{k_{1}+1}\otimes\ldots\otimes a_{k_{1}+n_{1}})\otimes$
$\ldots\otimes a_{k_{2}}\otimes g_{2}(a_{k_{2}+1}\otimes a_{k_{2}+n_{2}+1})\otimes\ldots$
$\otimes a_{k_{i}}\otimes g_{i}(a_{k_{i}+1}\otimes\ldots\otimes a_{k_{i}+n_{i}})\otimes\ldots\otimes a_{n+n_{1}+\ldots+n_{i}-i},$
where ${\varepsilon}\coloneqq\sum_{p=1}^{i}n_{p}k_{p}$. The brace bracket $f\{g\}$ is the Gerstenhaber circle product $f\circ g$ [14]. One can check directly [38, (4+5)]
(4.4) $x\cdot y=(-1)^{|x|}m_{2}\{x,y\},$
(4.5) $\delta x=m_{2}\circ x-(-1)^{|x|+1}x\circ m_{2}.$
The Gerstenhaber bracket on $C^{*,*}(H^{*},H^{*})$ is defined as follows
(4.6) $a,b]\coloneqq a\circ b-(-1)^{(|a|+1)(|b|+1)}b\circ a.$
Then $\big(C^{*,*}(H^{*},H^{*}),\delta,[-,-]\big)$ is a Differential Graded Lie Algebra (DGLA).
In [23, Section 4.3] Kadeishvili showed that a twisting element $m\coloneqq m_{3}+m_{4}+\ldots\in C^{*,2-*}(H^{*},H^{*})$, $m_{p}\in C^{p,2-p}(H^{*},H^{*})$, defines a minimal $C_{\infty}$-enrichment of $H^{*}$ iff $\delta m=m\circ m$. Equivalently
(4.7) $\delta m=-\frac{1}{2}[m,m].$
Furthermore, the action of a $C_{\infty}$-isotopy $\mathrm{Id}+\phi\coloneqq\mathrm{Id}+\phi_{2}+\ldots$ where $\phi_{i}\in C^{i,1-i}(H^{*},H^{*})$, $i\geq 2$ on the set $Tw\big(C^{*,*}(H^{*})\big)$ of all twisting elements can be described in terms of $\delta$, product $\cdot$, and brace operations recursively as follows [23, (19)]
[…]
Corollary 4.1 (Gauge equivalence).
The equivalence of twisting cochains defined in (4.8) coincides with the usual gauge equivalence of Maurer–Cartan (MC) elements in the DGLA $\mathfrak{g}_{{H^{*}}}\coloneqq\big(C^{*,*}(H^{*},H^{*}),\delta,[-,-]\big)$:
(4.9) $m^{\prime}\sim m\quad\Longleftrightarrow\quad m^{\prime}=e^{{\rm ad}_{p}}\ast m\coloneqq e^{{\rm ad}_{p}}(m)-\frac{e^{{\rm ad}_{p}-1}}{{\rm ad}_{p}}(\delta p)$
for some degree zero $p=\phi_{2}+\phi_{3}+\ldots\in\mathfrak{g}_{H^{*}}^{0}$ and
${\rm ad}_{p}(x)\coloneqq[p,x].$
Here the degree $|\phi|$ of $\phi\in C^{i,j}(H^{*},H^{*})$ is defined by $|\phi|\coloneqq(i+j-1)$.
While Kadeishvili’s original derivation in [23] establishes the structural action of $C_{\infty}$-isotopies, the explicit identification with the Maurer-Cartan gauge action (Corollary 4.1) follows from the Voronov-Gerstenhaber work [38] and Goldman-Millson work [15, Lemma 2.8]. Specifically, the brace operations $m\{p,\dots,p\}$ represent the higher-order terms of the exponential gauge action, and for $n=3,4$, these match the component-wise expansion of the $C_{\infty}$-morphism equation (4.1).
Next we shall prove the following useful Lemma.
Lemma 4.2 (Distributive Formula for the Gauge Action).
Let $\mathfrak{g}$} be a DGLA. For any MC element $m\in{\mathfrak{g}}$ and any $\tau\in{\mathfrak{g}}$, we have
(4.10) $e^{\text{ad}_{q}}\ast(m+\tau)=(e^{\text{ad}_{q}}\ast m)+e^{\text{ad}_{q}}(\tau).$
Proof.
By (4.9), we have
$e^{\text{ad}_{q}}\ast(m+\tau)=e^{\text{ad}_{q}}(m+\tau)-\frac{e^{\text{ad}_{q}}-1}{\text{ad}_{q}}(\delta q).$
We expand the first term in the RHS of the above equality:
$e^{\text{ad}_{q}}\ast(m+\tau)=e^{\text{ad}_{q}}(m)+e^{\text{ad}_{q}}(\tau)-\frac{e^{\text{ad}_{q}}-1}{\text{ad}_{q}}(\delta q).$
Now, regrouping the terms involving $m$ and $q$:
$e^{\text{ad}_{q}}\ast(m+\tau)=\underbrace{\left(e^{\text{ad}_{q}}(m)-\frac{e^{\text{ad}_{q}}-1}{\text{ad}_{q}}(\delta q)\right)}_{e^{\text{ad}_{q}}\ast m}+e^{\text{ad}_{q}}(\tau)$
which is Equality (4.10). ∎
(iii) For a Poincaré GCA $H^{*}$, let $CC^{*}(H^{*})$ denote the complex of cyclic multilinear forms on $H^{*}$. Given a cyclic Harrison cochain $f\in CC^{k,m}(H^{*},H^{*})$ consisting of all Harrison cochains that satisfy the cyclicity (3.3) we denote its associated $(k+1)$-form by $Tf$, defined by [29, Chapter 2]:
(4.11) $Tf(x_{1},\dots,x_{k},x_{k+1}):=\langle f(x_{1},\dots,x_{k}),x_{k+1}\rangle.$
The resulting cyclic cohomology is denoted by $HC^{*,*}(H^{*})$. (We use Loday’s algebraic cyclic cohomology in the sense of Connes–Loday for graded commutative algebras with invariant pairing.) It follows from Hamilton–Lazarev’s works [18, Theorems 3.8], [19, Corollary 8.12] that
(4.12) $HC^{(i+1),*}(H^{*})=Harr^{i,*}(H^{*},H^{*})\text{ if }i\geq 3.$
4.2. Isotopy modulo $k$ for minimal $C_{\infty}$-algebra enhancements
Definition 4.3 (Equality Modulo k).
Let $A_{1}=(H^{*},\{m_{i}^{(1)}\})$ and $A_{2}=(H^{*},\{m_{i}^{(2)}\})$ be two minimal $C_{\infty}$-algebra enhancements of a GCA $H^{*}$ of finite type. We say that $A_{1}$ and $A_{2}$ are equal modulo $k$, written as $A_{1}=A_{2}\pmod{k}$, if their defining operations are identical up to order $k$. That is:
$m_{i}^{(1)}=m_{i}^{(2)}\quad\text{for all }i\leq k.$
Definition 4.4 (Isotopy Modulo k).
We define the relation “isotopy modulo kk”, denoted by $\sim_{k}$, on the set of minimal $C_{\infty}$-algebra enhancements of a GCA $H^{*}$ of finite type as follows. We say that $A_{1}$ is isotopic to $A_{2}$ modulo $k$, written as $A_{1}\sim_{k}A_{2}$, if there exists a $C_{\infty}$-isotopy $\phi$ such that the transformed algebra $\phi(A_{1})$ is equal to $A_{2}$ modulo $k$.
Proposition 4.5.
The isotopy modulo $k$ is an equivalence relation for any integer $k\geq 2$.
Remark 4.6 (Filtered pronilpotent DGLA).
We note that the notion of isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, on the set of minimal $C_{\infty}$-algebra enhancements can be described using a filtration on $g_{H^{*}}$ that have been considered by Hamilton-Lazarev in [18]. Namely we consider the following filtration
(4.13) $F^{k}{\mathfrak{g}}_{H^{*}}\coloneqq\prod_{n\geq k}C^{n,*}(H^{*},H^{*})$
consists of Harrison cochains of arity $\geq k$. Then we have
${\mathfrak{g}}_{H^{*}}=F^{1}{\mathfrak{g}}_{H^{*}}\supset F^{2}{\mathfrak{g}}_{H^{*}}\supset\ldots$
such that
1. $\delta(F^{p}{\mathfrak{g}}_{H^{*}})\subset F^{p}{\mathfrak{g}}_{H^{*}}$,
2. $F^{p}{\mathfrak{g}}_{H^{*}},F^{q}{\mathfrak{g}}_{H^{*}}]\subset F^{p+q}{\mathfrak{g}}_{H^{*}}$,
3. $\bigcap_{p}F^{p}{\mathfrak{g}}_{H^{*}}=0.$
Furthermore, two elements $m,m^{\prime}\in{\mathfrak{g}}_{H^{*}}$ are isotopic modulo $k$ iff there exists $p\in{\mathfrak{g}}^{0}_{H^{*}}$ such that $e^{{\rm ad}_{p}}\ast(m)=m^{\prime}\mod F^{k}{\mathfrak{g}}_{H^{*}}$. In this filtered DGLA formalism, the relation $\sim_{k}$ is effectively a group action on the quotient ${\mathfrak{g}}_{H^{*}}/F^{k}{\mathfrak{g}}_{H^{*}}$. Hence, Proposition 4.5 is immediate. Note that Hamilton-Lazarev considered minimal $C_{\infty}$-algebra structures modulo filtration and obstructions to extendability of a minimal $C_{\infty}$-algebra structure modulo $F^{n}{\mathfrak{g}}_{H^{*}}$ to a minimal $C_{\infty}$-algebra structure modulo $F^{n+1}{\mathfrak{g}}_{H^{*}}$.
If $H^{*}$ is a PGCA then we denote by ${\mathfrak{g}}_{H^{*},\,cycl,\,unit}$ the sub-DGLA of ${\mathfrak{g}}_{H^{*}}$ consisting of all normalized cyclic Harrison cochains. We abbreviate $F^{k}{\mathfrak{g}}$ as $F^{k}$ if no confusion can occur.
Theorem 4.7.
Let $A_{1}$ and $A_{2}$ be two minimal $C_{\infty}$-algebra enhancements of a finite type graded commutative algebra $H^{*}$ over a field $\mathbb{F}$} of characteristic 0.
a) The following are equivalent:
  1. (1)
    $A_{1}$ and $A_{2}$ are isotopic as minimal $C_{\infty}$-algebras.
  2. (2)
    $A_{1}$ and $A_{2}$ are isotopic modulo $k$ for all integers $k\geq 3$.
b) If moreover $H^{*}$ is a PGCA and $A_{1},A_{2}$ are minimal unital cyclic $C_{\infty}$-algebra enhancements, then the following are equivalent:
  1. (1)
    $A_{1}$ and $A_{2}$ are isotopic as minimal $C_{\infty}$-algebras.
  2. (2)
    $A_{1}$ and $A_{2}$ are isotopic as minimal unital cyclic $C_{\infty}$-algebras, i.e, there exists $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*},\,cyc,\,unit}$ such that $A_{2}=e^{{\rm ad}_{p}}\ast A_{1}$.
  3. (3)
    $A_{1}$ and $A_{2}$ are isotopic modulo $k$ as minimal unital cyclic $C_{\infty}$-algebras, for all integers $k\geq 3$.
Proof.
a) Let us prove the equivalence in the first assertion. Minimal $C_{\infty}$-algebra enhancements of $H^{*}$ correspond to Maurer–Cartan elements in the filtered Harrison DGLA ${\mathfrak{g}}_{H^{*}}$. Isotopy modulo $k$ is precisely gauge equivalence in the quotient DGLA ${\mathfrak{g}}_{H^{*}}/F^{k+1}$. Since $H^{*}$ is of finite type $\mathfrak{g}$ is complete with respect to the arity filtration
[…]
Thus two Maurer–Cartan elements are gauge equivalent if and only if they are equivalent in every quotient ${\mathfrak{g}}_{H^{*}}/F^{k+1}$. This proves $(2)\Rightarrow(1)$. The other implications are immediate.
b) Let us prove the second assertion. The equivalence (1) $\iff$ (2) is established by Hamilton-Lazarev [18, Theorems 5.5, 5.10].
The equivalence (2) $\iff$ (3) holds because ${\mathfrak{g}}_{H^{*},\,cycl,\,unit}$ is a filtered sub-DGLA of the filtered DGLA ${\mathfrak{g}}_{H^{*}}$. ∎
4.3. The Primary Obstruction: Equivalence Modulo 3
The first step in comparing two $C_{\infty}$-algebra enhancements is to analyze their isotopy modulo $3$. The morphism equation at this level reveals the first potential obstruction to isotopy modulo $4$, which lives in Harrison cohomology. The following proposition shows that this primary obstruction completely characterizes isotopy modulo 3.
Proposition 4.8 (Primary obstruction).
Let $H^{*}$ be a GCA of finite type. Two minimal $C_{\infty}$-algebra enhancements, $A_{1}=(H^{*},\{m_{i}^{(1)}\})$ and $A_{2}=(H^{*},\{m_{i}^{(2)}\})$, are isotopic modulo 3 if and only if their primary obstruction classes coincide in Harrison cohomology:
(4.14) $A_{1}\sim_{3}A_{2}\iff[m_{3}^{(1)}]=[m_{3}^{(2)}]\in\mathrm{Harr}^{3,-1}(H^{*},H^{*}).$
If, moreover $H^{*}$ is PGCA and $A_{1}$ and $A_{2}$ are unital and cyclic, then
(4.15) $A_{1}\sim_{3}A_{2}\iff[Tm_{3}^{(1)}]=[Tm_{3}^{(2)}]\in HC^{4,-1}(H^{*}).$
Proof.
Let us first prove the equivalence relation (4.14).
(i) ($\implies$) The “only if” statement. Assume $A_{1}\sim_{3}A_{2}$. By definition, this implies the existence of a $C_{\infty}$-isotopy $\phi=(\mathrm{Id},\phi_{2},\ldots)$ such that the transformed algebra $A^{\prime\prime}_{1}=\phi(A_{1})$ satisfies $m_{3}^{(1^{\prime\prime})}=m_{3}^{(2)}$. The transformation formula from Equation (4.1) for $n=3$ simplifies to $m_{3}^{(1^{\prime\prime})}=m_{3}^{(1)}-\delta\phi_{2}$. Substituting the equivalence condition, we get $m_{3}^{(2)}=m_{3}^{(1)}-\delta\phi_{2}$. Rearranging gives:
$m_{3}^{(1)}-m_{3}^{(2)}=\delta\phi_{2}=d_{Harr}\phi_{2}.$
This shows that the cochain $(m_{3}^{(1)}-m_{3}^{(2)})$ is a coboundary, meaning their Harrison cohomology classes are equal.
[…]
$m_{3}^{(1)}-m_{3}^{(2)}=\delta\phi_{2}.$
We use this $\phi_{2}$ to construct a $C_{\infty}$-isotopy. Let $\phi=(\phi_{1}=\mathrm{Id},\phi_{2},\phi_{3}=0,\ldots)$. Let $A^{\prime\prime}_{1}=\phi(A_{1})$ be the transformed algebra. We check if $A^{\prime\prime}_{1}=A_{2}\pmod{3}$. The new $m_{2}$ operation, $m_{2}^{(1^{\prime\prime})}$, is equal to $m_{2}^{(1)}$, and since both are minimal $C_{\infty}$-algebra enhancements of $H^{*}$, we have $m_{2}^{(1)}=m_{2}^{(2)}$. For the $m_{3}$ operation, the transformation formula gives:
(4.16) $m_{3}^{(1^{\prime\prime})}=m_{3}^{(1)}-\delta\phi_{2}.$
Substituting our expression for $\delta\phi_{2}$, we find:
$m_{3}^{(1^{\prime\prime})}=m_{3}^{(1)}-(m_{3}^{(1)}-m_{3}^{(2)})=m_{3}^{(2)}.$
Thus, we have constructed a $C_{\infty}$-isotopy $\phi$ such that $\phi(A_{1})=A_{2}\pmod{3}$. Therefore, $A_{1}\sim_{3}A_{2}$.
Finally, the equivalence relation (4.15) follows from (4.14) and (4.12).
4.4. The higher obstructions: Isotopy Modulo $k\geq 4$
Let us begin with the secondary obstruction for extending isotopy modulo 3 to isotopy modulo 4. Assume that $m\sim_{3}m^{\prime}$. Then there exists an element $p=\phi_{2}+\ldots\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such $e^{{\rm ad}_{p^{\prime}}}\ast m=m\mod F^{3}{\mathfrak{g}}_{H^{*}}$. By (4.16), we have
(4.17) $m_{3}^{\prime}=m_{3}-\delta\phi_{2}.$
We are interested in the question if there exists an element $p=(\phi_{2},\phi_{3},\ldots)\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such that
(4.18) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{4}{\mathfrak{g}}_{H^{*}}.$
In this case, we have
(4.19) $m_{4}^{\prime}=m_{4}+[\phi_{2},m_{3}]-\delta\phi_{3}-\frac{1}{2}[\phi_{2},\delta\phi_{2}].$
Lemma 4.9.
The cochain
(4.20) $\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4}^{\prime})\coloneqq m_{4}-m_{4}^{\prime}+[\phi_{2},m_{3}]-\frac{1}{2}[\phi_{2},\delta\phi_{2}]$
is a cocycle in the Harrison complex $Harr^{4,-2}(H^{*},H^{*})$. Furthermore, the cohomology class of $\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4}^{\prime})$ can be represented by $m_{4}-m_{4}^{\prime}+[\phi_{2},m_{3}]$.
Proof.
From the equation $\delta m=-\frac{1}{2}[m,m]$ we obtain
$\delta m_{4}=-\frac{1}{2}[m_{3},m_{3}].$
Similarly, $\delta m_{4}^{\prime}=-\frac{1}{2}[m_{3}^{\prime},m_{3}^{\prime}]$. Recall that $m_{3}^{\prime}=m_{3}-\delta\phi_{2}$. Hence
(4.21) $\delta(m_{4}-m_{4}^{\prime})=-[m_{3},\delta\phi_{2}]+\frac{1}{2}[\delta\phi_{2},\delta\phi_{2}]$
It follows
$\delta\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4})=-[m_{3},\delta\phi_{2}]+\frac{1}{2}[\delta\phi_{2},\delta\phi_{2}]+\delta[\phi_{2},m_{3}]-\frac{1}{2}\delta[\phi_{2},\delta\phi_{2}]=0.$
Thus, the last assertion of Lemma 4.9 follows from the observation that $2[\phi_{2},\delta\phi_{2}]=\delta[\phi_{2},\phi_{2}]$. ∎
It follows from Lemma 4.9 that the existence of an element $p=(\phi_{2},\phi_{3},\ldots)\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such that Equation (4.18) holds is equivalent to the vanishing of the cohomology class of $(m_{4}-m_{4}^{\prime}+[\phi_{2},m_{3}])$ in $Harr^{4,-2}(H^{*},H^{*})$.
Lemma 4.10.
Assume that $m=(m_{3},m_{4},\ldots)$ and $m^{\prime}=(m_{3}^{\prime},m_{4}^{\prime},\ldots)$ are given two minimal $C_{\infty}$-algebra enhancements of a finite type GCA $H^{*}$. Let $\phi^{\prime}\in C^{2,-1}(H^{*},H^{*})$ satisfy the condition $\delta\phi_{2}^{\prime}=m_{3}-m_{3}^{\prime}$, i.e., $\phi_{2}^{\prime}=\phi_{2}+f$ for some $f\in C^{2,-1}(H^{*},H^{*})$ such that $\delta f=0$. Then we have
(4.22) $\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2}+f,m_{4}^{\prime})=\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4}^{\prime})+\frac{1}{2}[f,m_{3}+m_{3}^{\prime}].$
Furthermore, the cohomology class $\frac{1}{2}[f,m_{3}+m_{3}^{\prime}]=[f,m_{3}]$ depends only on $f]\in Harr^{2,-1}(H^{*},H^{*})$.
Proposition 4.11 (The secondary obstruction space).
Let $m=(m_{3},m_{4},\ldots,)$ and $m^{\prime}=(m_{3},m_{4},\ldots)$ be two minimal $C_{\infty}$-algebra enhancements of a finite type GCA $H^{*}$. Assume that $m\sim_{3}m^{\prime}$.
1) A necessary and sufficient condition for $m\sim_{4}m^{\prime}$ is that the secondary obstruction affine space
${\mathcal{K}}_{4}(m_{3},m_{4},m_{3}^{\prime},m_{4}^{\prime})\coloneqq\{[\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4}^{\prime})]\in Harr^{4,-2}(H^{*},H^{*}):\delta\phi_{2}=m_{3}-m_{3}^{\prime}\}.$
contains the zero in $Harr^{4,-2}(H^{*},H^{*})$.
2) Equivalently, $m\sim_{4}m^{\prime}$ if and only if for some $\phi\in C^{2,-1}(H^{*},H^{*})$ such that $\delta\phi_{2}=m_{3}-m_{3}^{\prime}$ we have
(4.23) $\tilde{\kappa}_{4}(m_{3},m_{4},\phi_{2},m_{4}^{\prime})]\in{\rm ad}_{[m_{3}]}(Harr^{2,-1}(H^{*},H^{*})).$
3) Equivalently, $m\sim_{4}m^{\prime}$ if and only if we have
(4.24) ${\mathcal{K}}_{4}(m_{3},m_{4},m_{3}^{\prime},m_{4}^{\prime})\subset{\rm ad}_{[m_{3}]}(Harr^{2,-1}(H^{*},H^{*})).$
We shall derive Lemma 4.10 and Proposition 4.11 from Theorem 4.12 below. Denote by
$\mathrm{Proj}_{k}:{\mathfrak{g}}_{H^{*}}\to F^{k}{\mathfrak{g}}_{H^{*}}/F^{k+1}{\mathfrak{g}}_{H^{*}}$
the natural projection.
Theorem 4.12 (Higher obstruction classes).
Assume that $m\sim_{k-1}m^{\prime}$, $k\geq 4$, are two minimal $C_{\infty}$-algebra enhancements of a finite type GCA $H^{*}$. Set
$\widetilde{\mathcal{K}}_{k}(m,m^{\prime})\coloneqq$ $\{\tilde{\kappa}_{k}(p,m,m^{\prime})\coloneqq\mathrm{Proj}_{k}(e^{{\rm ad}_{p}}\ast m-m^{\prime})\in F^{k}{\mathfrak{g}}_{H^{*}}/F^{k+1}{\mathfrak{g}}_{H^{*}},$
(4.25) $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*}},\,e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}{\mathfrak{g}}_{H^{*}}\}.$
1) Then $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ consists of cocycles in $C^{k,(2-k)}(H^{*},H^{*})$ and the image ${\mathcal{K}}_{k}(m,m^{\prime})$ of $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ in $Harr^{k,-(k-2)}(H^{*},H^{*})$ depends only on $m\mod F^{k}=(m_{3},\ldots,m_{k})$ and $m\mod F^{k}=(m_{3}^{\prime},\ldots,m_{k}^{\prime})$.
2) Furthermore
(4.26) ${\mathcal{K}}_{k}(m,m^{\prime})\ni 0\Longleftrightarrow m\sim_{k}m^{\prime}.$
3) The affine space ${\mathcal{K}}_{k}(m,m^{\prime})$ is modeled on the linear space ${\mathcal{K}}_{k}(m^{\prime},m^{\prime})$, i.e., there exists a free transitive action of ${\mathcal{K}}_{k}(m^{\prime},m^{\prime})$ on ${\mathcal{K}}_{k}(m,m^{\prime})$.
4) A necessary condition for ${\mathcal{K}}_{k}(m,m^{\prime})\ni 0$ is
(4.27) ${\mathcal{K}}_{k}(m,m)={\mathcal{K}}_{k}(m^{\prime},m^{\prime}).$
We shall provide a proof of Theorem 4.12 which works for any filtered DGLA $({\mathfrak{g}},\delta,[-,-],F^{k})_{k\in{\mathbb{N}}^{+}}$. So we shall write $\mathfrak{g}$} instead of ${\mathfrak{g}}_{H^{*}}$ and regard minimal $C_{\infty}$-algebra enhancements $m$ as MC elements. Suppose that $m,m^{\prime}\in F^{3}{\mathfrak{g}}^{1}$ are two MC elements such that $m\sim_{k-1}m^{\prime}$, $k\geq 3,$ and there exists an element $p\in F^{2}{\mathfrak{g}}^{0}$ with the following property (cf. (4.25))
(4.28) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}{\mathfrak{g}}.$
Denote by
$\mathrm{Proj}_{k}:{\mathfrak{g}}\to F^{k}{\mathfrak{g}}/F^{k+1}{\mathfrak{g}}$
the natural projection. The $k$-the obstruction class $\tilde{\kappa}_{k}(p,m,m^{\prime})]$ is the cohomology class in $H^{*}(F^{k}/F^{k+1},\delta)$ that measure the failure to extend this equivalence to the next level, namely (cf. (4.25)):
(4.29) $\tilde{\kappa}_{k}(p,m,m^{\prime})\coloneqq\mathrm{Proj}_{k}(e^{{\rm ad}_{p}}\ast m-m^{\prime})\in F^{k}/F^{k+1}.$
We shall prove Lemmas 4.13, 4.14, 4.15 and Proposition 4.16 before completing the proof of Theorem 4.12.
Lemma 4.13 (Closedness).
$\tilde{\kappa}_{k}(p,m,m^{\prime})$ is cocycle in ${\mathfrak{g}}/F^{k+1}$.
Proof of Lemma 4.13.
Set $M\coloneqq e^{{\rm ad}_{p}}\ast m$. Then $M$ is also a MC element, and hence
$\delta M+\frac{1}{2}[M,M]=0.$
By the assumption (4.28), we have
(4.30) $M=m+\tilde{\kappa}_{k}+R\text{ where }R\in F^{k+1}.$
Substitute $M=m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})+R$ into the MC equation for $M$ and work modulo $F^{k+1}$:
$\delta\big(m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big)+\frac{1}{2}\big[m^{\prime}+\tilde{\kappa}_{k}(p,m,^{\prime}),m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
It follows that
$\left(\delta m^{\prime}+\frac{1}{2}[m^{\prime},m^{\prime}]\right)+\delta\tilde{\kappa}_{k}(p,m,m^{\prime})+\big[m^{\prime},\tilde{\kappa}_{k}(p,m,m^{\prime})\big]$
$\frac{1}{2}\big[\tilde{\kappa}_{k}(p,m,m^{\prime}),\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
The first term in the above equality vanishes since $m^{\prime}$ is a MC element. The last two terms vanish since $m^{\prime}\in F^{3}$ and $\tilde{\kappa}_{k}\in F^{k}$, $k\geq 3$. It follows that $\delta\tilde{\kappa}_{k}(p,m,m^{\prime})=0\mod F^{k+1}$. This completes the proof of Lemma 4.13. ∎
Lemma 4.14 (Transitivity).
If $p,\bar{p}\in F^{2}{\mathfrak{g}}^{0}$ satisfy
(4.31) $e^{{\rm ad}_{p}}\ast m=e^{{\rm ad}_{\bar{p}}}\ast m=m^{\prime}\mod F^{k-1}$
then
(4.32) $\big[\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=[\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})]+[\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})]$
where $q\in F^{2}{\mathfrak{g}}^{0}$ such that
(4.33) $e^{{\rm ad}_{q}}=e^{{\rm ad}_{\bar{p}}}\cdot(e^{{\rm ad}_{p}})^{-1}.$
Proof of Lemma 4.14.
Step 1: Defining $q$ from Equation (4.33) and proving that $q$ is a stabilizer of $m^{\prime}$ modulo $F^{k-1}$.
Let $p,\bar{p}\in F^{2}\mathfrak{g}^{0}$ satisfy (4.31). Then we have
(4.34) $X\coloneqq e^{\text{ad}_{p}}\ast m=m^{\prime}+\alpha,\,\alpha\in F^{k},$
(4.35) $\mathrm{Proj}_{k}(\alpha)=\tilde{\kappa}_{k}(p,m,m^{\prime}),$
(4.36) $Y\coloneqq e^{\text{ad}_{\bar{p}}}\ast m=m^{\prime}+\beta,\,\beta\in F^{k},$
(4.37) $\mathrm{Proj}_{k}(\beta)=\tilde{\kappa}_{k}(\bar{p},m,m^{\prime}).$
Now, define $q\in F^{2}\mathfrak{g}^{0}$ via the group structure of the gauge group:
$e^{\text{ad}_{q}}=e^{\text{ad}_{\bar{p}}}\cdot(e^{\text{ad}_{p}})^{-1}$
by the BCH formula:
$q=\bar{p}-p+\frac{1}{2}[\bar{p},-p]+\frac{1}{12}\left([\bar{p},[\bar{p},-p]]+[-p,[-p,\bar{p}]]\right)-\frac{1}{24}[-p,[\bar{p},[\bar{p},-p]]]+\dots$
$(\bar{p}-p)-\frac{1}{2}[\bar{p},p]+\frac{1}{12}\left([\bar{p},[\bar{p},p]]+[p,[p,\bar{p}]]\right)-\frac{1}{24}[p,[\bar{p},[\bar{p},p]]]+\dots.$
Since $p,\bar{p}\in F^{2}\mathfrak{g}^{0}$ every term in the BCH expansion involves at least one $p$ or $\bar{p}$, so $q\in F^{2}$.
From (4.33), (4.34), and (4.36), we obtain
$e^{{\rm ad}_{q}}\ast X=Y.$
Substitute $X=m^{\prime}+\alpha$, we have
(4.38) $e^{{\rm ad}_{q}}\ast(m^{\prime}+\alpha)=Y.$
Using the distributive formula (4.10) we derive from (4.38):
(4.39) $e^{{\rm ad}_{q}}\ast(m^{\prime}+\alpha)=(e^{{\rm ad}_{q}}\ast m^{\prime})+e^{{\rm ad}_{q}}(\alpha).$
Taking into account (4.38) and (4.36), we derive from (4.39) the following master equation:
(4.40) $(e^{{\rm ad}_{q}}\ast m^{\prime})+e^{{\rm ad}_{q}}(\alpha)=m^{\prime}+\beta.$
We know $\alpha,\beta\in F^{k}$. Let’s expand $e^{\text{ad}_{q}}(\alpha)$:
$e^{\text{ad}_{q}}(\alpha)=\alpha+[q,\alpha]+\dots$
Since $q\in F^{2}$ and $\alpha\in F^{k}$, the bracket $q,\alpha]$ is in $F^{k+1}$. Therefore:
$e^{\text{ad}_{q}}(\alpha)\equiv\alpha\pmod{F^{k+1}}$.
Now substitute this back into the master equation (4.40)
$(e^{\text{ad}_{q}}\ast m^{\prime})+\alpha\equiv m^{\prime}+\beta\mod{F^{k+1}}.$
Rearrange to isolate the gauge action on $m^{\prime}$:
(4.41) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\equiv\beta-\alpha\mod{F^{k+1}}.$
Because $\alpha$ and $\beta$ are both in $F^{k}$, their difference $(\beta-\alpha)$ is also in $F^{k}$. Hence, (4.41) implies:
(4.42) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\in F^{k},$
i.e, $q$ is a stabilizer of $m^{\prime}$ modulo $F^{k-1}$. $\Box$
Step 2. The final projection to $F^{k}$.
From (4.42) and we have
$\text{Proj}_{k}(e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime})=\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime}).$
Taking into account (4.35) and (4.37):
$\text{Proj}_{k}(\beta)=\tilde{\kappa}_{k}(\bar{p},m,m^{\prime}),$
$\text{Proj}_{k}(\alpha)=\tilde{\kappa}_{k}(p,m,m^{\prime}),$
Equation (4.41) gives the final, clean identity:
$\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})=\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})-\tilde{\kappa}_{k}(p,m,m^{\prime}).$
Equivalently,
$\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})]=[\tilde{\kappa}_{k}(p,m,m^{\prime})]+[\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})].$
This completes the proof of Lemma 4.14. ∎
Lemma 4.15 (Existence of a free action).
Let $p\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}$. Set
$\tilde{\kappa}_{k}(p,m,m^{\prime})\coloneqq\mathrm{Proj}_{k}(e^{{\rm ad}_{p}}\ast m-m^{\prime})\in\tilde{\mathcal{K}}_{k}(m,m^{\prime}).$
Let $q\in F^{2}{\mathfrak{g}}^{0}$ such that
$\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})]\in\mathcal{K}_{k}(m^{\prime},m^{\prime}).$
Then there exists $\bar{p}\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{\bar{p}}}\ast m=m^{\prime}\pmod{F^{k-1}}$ and
(4.43) $\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})]=[\tilde{\kappa}_{k}(p,m,m^{\prime})]+[\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})].$
Proof of Lemma 4.15.
First we define $\bar{p}$ via the group composition in the gauge group (cf. (4.33)):
(4.44) $e^{\text{ad}_{\bar{p}}}=e^{\text{ad}_{q}}\cdot e^{\text{ad}_{p}}.$
Since $p$ and $q$ are both in $F^{2}\mathfrak{g}^{0}$, their BCH composition $\bar{p}$ is also in $F^{2}\mathfrak{g}^{0}$. From (4.44) we obtain:
(4.45) $e^{\text{ad}_{\bar{p}}}\ast m=e^{\text{ad}_{q}}\ast(e^{\text{ad}_{p}}\ast m).$
By our assumption on $p$ we have
(4.46) $e^{\text{ad}_{p}}\ast m=m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p},\quad R_{p}\in F^{k+1}.$
Applying the distributive formula (4.10) to (4.46), we obtain
(4.47) $e^{\text{ad}_{\bar{p}}}\ast m=(e^{\text{ad}_{q}}\ast m^{\prime})+e^{\text{ad}_{q}}\big(\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p}\big).$
Next we expand the stabilizer term. By our assumption on $q$:
(4.48) $e^{\text{ad}_{q}}\ast m^{\prime}=m^{\prime}+\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})+R_{q},\quad R_{q}\in F^{k+1}.$
We substitute the expansion from (4.48) into (4.47):
$e^{{\rm ad}_{\bar{p}}}\ast m=m^{\prime}+\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})+R_{q}+e^{{\rm ad}_{q}}\big(\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p}\big).$
Equivalently.
(4.49) $e^{{\rm ad}_{\bar{p}}}\ast m-m^{\prime}=\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})+R_{q}+e^{{\rm ad}_{q}}\big(\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p}\big).$
Now, project the LHS and RHS of Equation (4.49) to the $F^{k}$. Since $q\in F^{2}$ and $\tilde{\kappa}_{k}(p)\in F^{k}$, the adjoint term $e^{\text{ad}_{q}}(\tilde{\kappa}_{k}(p))=\tilde{\kappa}_{k}(p)+[q,\tilde{\kappa}_{k}(p)]+\dots$ satisfies:
(4.50) $\mathrm{Proj}_{k}\Big(e^{\text{ad}_{q}}\big(\tilde{\kappa}_{k}(p,m,m^{\prime})\big)\Big)=\tilde{\kappa}_{k}(p,m,m^{\prime}).$
because the brackets land in $F^{k+1}$. From (4.49) and (4.50) we obtain:
$\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})=\tilde{\kappa}_{k}(p,m,m^{\prime})+\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime}).$
Passing to cohomology classes:
$\tilde{\kappa}_{k}(\bar{p},m,m^{\prime})]=[\tilde{\kappa}_{k}(p,m,m^{\prime})]+[\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime})].$
This completes the proof of Lemma 4.15. ∎
Proposition 4.16 (Generalized Additivity and Adjointness).
Let $k\geq 4$. Assume that MC elements $m,m^{\prime}\in F^{3}{\mathfrak{g}}^{1}$ and $m\sim_{k-1}m^{\prime}$.
1) Assume that $m^{\prime\prime}\in F^{3}{\mathfrak{g}}^{1}$ and $m^{\prime\prime}\sim_{k-1}m^{\prime}$. Then for any $p,q\in F^{2}{\mathfrak{g}}_{0}$ such that
(4.51) $e^{{\rm ad}_{p}}\ast m-m^{\prime}$ $=0\mod F^{k-1},$
(4.52) $\tilde{\kappa}_{k}(p,m,m^{\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{p}}\ast m-m^{\prime}\big),$
(4.53) $e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}$ $=0\mod F^{k-1},$
(4.54) $\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}\big),$
the element $q\ast p\coloneqq BCH(q,p)$ defined by
(4.55) $e^{{\rm ad}_{q\ast p}}=e^{{\rm ad}_{q}}\cdot e^{{\rm ad}_{p}}$
satisfies the following equations:
(4.56) $e^{q\ast p}\ast m$ $=m^{\prime\prime}\mod F^{k-1},$
(4.57) $\tilde{\kappa}_{k}(q\ast p,m,m^{\prime\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{q\ast p}}\ast m-m^{\prime\prime}\big)=\tilde{\kappa}(p,m,m^{\prime})+\tilde{\kappa}(q,m^{\prime},m^{\prime\prime}).$
2) There exist $p,q\in F^{2}{\mathfrak{g}}^{0}$ such that
(4.58) ${\mathcal{K}}_{k}(m,m^{\prime})=e^{{\rm ad}_{p}}\big({\mathcal{K}}_{k}(m^{\prime},m)\big),$
(4.59) ${\mathcal{K}}_{k}(m,m)=e^{{\rm ad}_{q}}\big({\mathcal{K}}_{k}(m^{\prime},m^{\prime})\big).$
3) The space
${\mathcal{K}}_{k}(p,m^{\prime},m^{\prime}):\big\{[\kappa_{k}(q,m^{\prime},m^{\prime})]\in Harr^{k,-(k-2)}(H^{*},H^{*}):$
$q\in F^{2}{\mathfrak{g}}_{0},\,e^{{\rm ad}_{q}}\ast m^{\prime}=m\mod F^{k}\big\}$
is a subgroup of the additive group $Harr^{k,-(k-2)}(H^{*},H^{*})$.
Proof of Proposition 4.16.
1) Note that the first assertion of Proposition 4.16is a generalization of Lemma 4.15 and can be proved in the same way. Let $m,m^{\prime},m^{\prime\prime}$ be as in the assumption of Proposition 4.16. For any $p\in F^{2}{\mathfrak{g}}_{0}$, we define the gauge defect $\Delta(p;m,m^{\prime})$ as follows:
[…]
Substitute $e^{\text{ad}_{p}}*m=m^{\prime}+\Delta(p;m,m^{\prime})$ into Equation
[…]
we obtain
(4.61) $\Delta(q*p;m,m^{\prime\prime})=e^{\text{ad}_{q}}*(m^{\prime}+\Delta(p;m,m^{\prime}))-m^{\prime\prime}.$
Applying the distributive formula (4.10) to the RHS of Equation (4.61), we derive
(4.62) $e^{\text{ad}_{q}}*(m^{\prime}+\Delta(p;m,m^{\prime}))=(e^{\text{ad}_{q}}*m^{\prime})+e^{\text{ad}_{q}}(\Delta(p;m,m^{\prime})).$
From (4.61) and (4.62), we obtain
(4.63) $\Delta(q*p;m,m^{\prime\prime})=\underbrace{(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})}_{\Delta(q;m^{\prime},m^{\prime\prime})}+e^{\text{ad}_{q}}(\Delta(p;m,m^{\prime})).$
Recalling (4.53),
$(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})=0\mod F^{k-1}.$
By (4.60)
$\Delta(p,m,m^{\prime})=\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p},\,R_{p}\in F^{k+1}.$
Since $q\in F^{2}$, we have $q,\Delta(p,m,m^{\prime})]\in F^{k+1}$. Hence
(4.64) $e^{\text{ad}_{q}}(\Delta(p;m,m^{\prime}))=\tilde{\kappa}_{k}(p,m,m^{\prime})+R_{p}^{\prime},\,R_{p}^{\prime}\in F^{k+1}.$
From (4.64), (4.63), taking into account (4.53), we conclude
$\Delta(p,m,m^{\prime})\in F^{k},$
which proves (4.56) in the first assertion of Proposition 4.16.
Taking into account (4.53), (4.54), we derive from (4.63) and (4.64):
$\tilde{\kappa}_{k}(q*p,m,m^{\prime\prime})=\tilde{\kappa}_{k}(q,m,m^{\prime})+\tilde{\kappa}_{k}(p,m,m^{\prime})\mod F^{k}.$
This is Equation (4.57) in the first assertion of Proposition 4.16.
(2) Proposition 4.16 (2) follows from Proposition 4.16(1).
(3) By Proposition 4.16(1), ${\mathcal{K}}_{k}(m,m)$ is a monoid. By Lemma 4.14, for any $x\in{\mathcal{K}}_{k}(m,m)$ there exists x-x. Hence, ${\mathcal{K}}_{k}(m,m)$ is a additive subgroup of $Harr^{k,(2-k)}(H^{*},H^{*})$. This completes the proof of Proposition 4.16. ∎
Proof of Theorem 4.12.
1) The first assertion of Theorem 4.12 follows from Lemma 4.13 and Formula (4.29), noting that $m,m^{\prime}\in F^{3}{\mathfrak{g}}^{1}$ and $p\in F^{2}{\mathfrak{g}}^{0}$.
2) The second assertion follows from (4.29) and the Formula of Gauge Action (4.9).
3) The third assertion of Theorem 4.12 follows from Lemmas 4.14, 4.15.
4) The last assertion follows the third one and Proposition 4.16. ∎
Proof of Lemma 4.10.
Lemma 4.10 follows directly from Theorem 4.12 and its proof. ∎
Proof of Proposition 4.11.
Proposition 4.11 follows directly from Theorem 4.12 and its proof. ∎
Let $r\geq 2$. Now, we are ready to classify real-and rational homotopy types of (r1)(r-1)-connected, smooth manifolds of dimension $n\leq l(r-1)+2$, $l\geq 4$.
Denote by $C^{*,*}_{cyl,\,unit}(H^{*},H^{*})$ the subcomplex of the Harrison cochain complex $C^{*,*}(H^{*},H^{*})$ consisting normalized cyclic Harrison cochains and denote by ${\mathfrak{g}}_{H^{*},\,cyc,\,unit}$ the corresponding sub-DGLA of ${\mathfrak{g}}_{H^{*}}$. Hamilton-Lazarev proved that for a Poincaré GCA $H$ there is 1-1 correspondence between the set $C_{\infty}(H^{*})$ of all minimal $C_{\infty}$-algebra enhancements of $H^{*}$ and the set $C_{\infty,\,cyc,\,unit}(H^{*})$ of all minimal unital cyclic $C_{\infty}$-algebra enhancements of $H^{*}$ [18, Theorems 5.5, 5.10].
Theorem 4.17.
Assume $M$ is a closed (r1)(r-1)-connected manifold of dimension $n\leq l(r-1)+2$, where $r\geq 2$ and $l\geq 4$. Then the rational homotopy type of $M$ is determined uniquely by the cohomology algebra $H^{*}\coloneqq H^{*}(M,{\mathbb{Q}})$ and the isotopy modulo (l2)(l-2) of the corresponding minimal unital cyclic $C_{\infty}$-algebra enhancements of $H^{*}$. In particular, if $n\leq 5r-3$ then the rational homotopy type of $M$ is defined by the corresponding cyclic cohomology in $HC^{4,-1}(H^{*})$ (i.e., the primary obstruction $Tm_{3}]$ vanishes), and if $n\leq 6r-4$ then two manifolds $M$ and $M^{\prime}$ have the same homotopy type if and only if they have the same cyclic cohomology in $HC^{4,-1}(H^{*})$, and the relative second obstruction class vanishes.
We begin with the following observation.
Lemma 4.18 (Degree restriction for normalized cyclic Harrison cochains).
Let $H^{*}$ be a (r1)(r-1)-connected PGCA of degree $n$ where $r\geq 2$, $n\leq l(r-1)+2$, and $l\geq 4$. For any set of inputs $\{a_{i}\}_{i=1}^{k}$ where $a_{i}\in H^{d_{i}}$, and a normalized cyclic Harrison cochain $\phi\in C^{k,(2-k)}_{cyl,\,unit}(H^{*},H^{*})$, the value $\phi(a_{1},\ldots,a_{k})$ can be nonzero only if $k\leq(l-2)$.
Proof.
Since $H^{*}$ is (r1)(r-1)-connected, $H^{k}=0$ for $1\leq k<r$. Consequently, the value of a normalized cyclic Harrison cochain $\phi\in C^{k,(2-k)}_{cycl,\,unit}(H^{*},H^{*})$ vanishes if any input of $\phi$ is the unit $\mathbf{1}\in H^{0}$. Thus, for any non-trivial contribution, we must have $d_{i}\geq r$ for all $i$.
Furthermore, a nonzero output $\phi(a_{1},\ldots,a_{k})$ must have a non-degenerate pairing with some element $a_{k+1}\in H^{d_{k+1}}(M)$. By Poincaré duality, the degrees must sum to the dimension $n$. Since $\phi\in C^{k,(2-k)}_{cycl,\,unit}(H^{*},H^{*})$ we have
$\Big(\sum_{i=1}^{k}d_{i}+2-k\Big)+d_{k+1}=n.$
Hence
(4.65) $\sum_{i=1}^{k+1}d_{i}=n+k-2\leq l(r-1)+k.$
Since $d_{i}\geq r$ we must have
(4.66) $\sum_{i=1}^{k+1}d_{i}\geq(k+1)r.$
From (4.65) and (4.66) we conclude
$(k+1)r\leq l(r-1)+k.$
It follows that
$(k+1)r-k\leq l(r-1)$
$\Longleftrightarrow(k+1)(r-1)\leq l(r-1)-1.$
Consequently, $k\leq l-2$.
Proof of Theorem 4.17.
Taking into account Lemma 4.18, we derive Theorem 4.17 immediately from Theorems 4.7 and 4.12. ∎
We also derive from Lemma 4.18 the following
Corollary 4.19.
Let $r\geq 2$ and let $\mathcal{A}$$^{\ast}$ be a (r1)(r-1)-connected Poincaré DGCA of degree $n$ admitting a Hodge homotopy. If $n\leq l(r-1)+2$, with $l\geq 4$, then $H^{*}({\mathcal{A}}^{*})$ carries a minimal unital cyclic $C_{\infty}$-algebra enhancement making it $C_{\infty}$-quasi-isomorphic to $\mathcal{A}$$^{\ast}$, and whose multiplications $m_{k}$ vanish for $k\geq l-1$. In particular, all the multiplications $m_{k}$ with $k\geq 5$ vanish if $n\leq 6r-4$ all the multiplications $m_{k}$ with $k\geq 4$ vanish if $n\leq 5r-3$. Also, all the multiplications $m_{k}$ with $k\geq 3$ vanish if $n\leq 4r-2$, and so $\mathcal{A}$$^{\ast}$ is formal in this case.
Corollary 4.19 is slightly better than [13, Theorem 2.12], which does not have “cyclic” in the formulation.
4.5. Example: A simply connected non-formal 8-manifold
In this subsection, we demonstrate our stratification of minimal $C_{\infty}$-algebra enhancements using the 8-dimensional simply connected symplectic manifold $\tilde{M}$ constructed by Fernández and Muñoz [12]. This manifold serves as a crucial benchmark for our theory, as it lies on the boundary of the dimension bound $n\leq 5r-2$ for $r=2$.
Recall that $\tilde{M}$ is obtained by a symplectic resolution of the orbifold $M/\mathbb{Z}_{3}$, where $M=\Gamma\backslash(H_{\mathbb{C}}\times\mathbb{C})$ is a compact nilmanifold. The Lie algebra $\mathfrak{g}$ of the complex Heisenberg group $H_{\mathbb{C}}$ has left-invariant complex 1-forms $\{\mu,\nu,\theta\}$ satisfying the structure equations:
(4.67) $d\mu=0,\quad d\nu=0,\quad d\theta=\mu\wedge\nu.$
The non-formality of the nilmanifold $M$ is well-known and is encoded in the non-triviality of the triple Massey product $\langle[\mu],[\nu],[\mu]\rangle$.
[…]
The failure of intrinsic formality in $\tilde{M}$ can be interpreted as follows:
  1. (1)
    Primary Obstruction: The operation $m_{3}$ corresponds to the triple Massey product. In the nilmanifold $M$, the relation $d\theta=\mu\wedge\nu$ forces a non-vanishing $m_{3}([\mu],[\nu],[\cdot])$ in the transferred structure.
  2. (2)
    Isotopy Modulo $k$: While $\tilde{M}$ and its formal counterpart have isomorphic cohomology rings (level $k=2$), they are not isotopic modulo $3$. The class $m_{3}]\in Harr^{3,-1}_{cyc,\,uni}(H^{*},H^{*})=H^{4,-1}(H)$ is non-zero, hence ${\mathcal{K}}_{3}(m,0)=[m_{3}]$ where $0=(m_{3}=0,\ldots,0)$ denotes the formal enhancement.
  3. (3)
[…]
(In general, we have the following observation. Let $\tilde{M}$ be the resolution of a nilmanifold $M/G$. If the primary obstruction $m_{3}$ of the invariant Nomizu model $((\Lambda\mathfrak{g}^{*})^{G},d)$ is non-trivial, and the resolution map $\pi^{*}$ is a $C_{\infty}$-morphism that is injective on the generators of $m_{3}$, then $\tilde{M}$ is non-formal.)
[…]
  • $x=[\mu\wedge\bar{\eta}+\bar{\mu}\wedge\eta]$
  • $y=[\nu\wedge\bar{\eta}+\bar{\nu}\wedge\eta]$
  • $z=[i\eta\wedge\bar{\eta}]$
where $\mu,\nu,\eta$ are the 1-forms of the complex Heisenberg group. The non-triviality of $m_{3}$ on these classes arises from the “hidden” variable $\theta$ in the relation $d\theta=\mu\wedge\nu$, which allows for the non-vanishing homotopy $H(x\wedge y)$.
Triple $(a,b,c)$ $m_{3}(a,b,c)$ Triple $(a,b,c)$ $m_{3}(a,b,c)$
$(x,x,y)$ $\mu\wedge\theta$ $(y,x,y)$ $2\nu\wedge\theta$
$(x,y,x)$ $2\mu\wedge\theta$ $(y,x,z)$ $\theta\wedge\eta$
$(x,y,y)$ $\nu\wedge\theta$ $(y,y,x)$ $\nu\wedge\theta$
$(x,y,z)$ $\theta\wedge\eta$ $(z,x,y)$ $\theta\wedge\eta$
$(y,x,x)$ $\mu\wedge\theta$ $(z,y,x)$ $\theta\wedge\eta$
Table 1. Non-vanishing $m_{3}$ operations in the invariant sector of $\tilde{M}$. computed via the Merkulov formula.
5. A new proof of Crowley–Nordström’s formality result
In this section, using the results in the previous sections, we give an alternative proof of the following result due to Crowley–Nordström, generalizing a previous result by Cavalcanti [5, Theorem 4], under the assumption that $\mathcal{A}$$^{\ast}$ admits a Hodge homotopy. We assume that the ground field $\mathbb{F}$$\mathbb{R}$}. [^1] This Section was written in collaboration with Domenico Fiorenza.
Theorem 5.1.
Let $\mathcal{A}$$^{\ast}$ be (r1)(r-1)-connected Poincaré DGCA of degree n=4r1n=4r-1 with $b^{r}\leq 3$ admitting a Hodge homotopy. Let $H^{*}\coloneqq H^{*}({\mathcal{A}})$. If there is an element $\varphi\in H^{2r-1}$ such that $\varphi\cdot-:H^{r}\to H^{3r-1}$ is an isomorphism, then $\mathcal{A}$$^{\ast}$ is intrinsically formal.
Proof of Theorem 5.1 for $b^{r}=1$.
By the degree reason, $m_{3}(x,y,z)$ vanishes unless $x,y,z\in H^{r}$. From (3.8) we have that $m_{3}(x,y,z)$ also vanishes when $x,y$, and $z$ lie on the same line in $H^{r}$. If $b^{r}=1$ this is necessarily the case, so $m_{3}$ identically vanishes and $\mathcal{A}$$^{\ast}$ is formal. ∎
5.1. The algebraic curvature tensor of $\mathcal{A}$$^{\ast}$
To prove Theorem 5.1 for $b^{r}=2$ and for $b^{r}=3$ we need some preparation. For the whole section, $({\mathcal{A}}^{\ast},d,d^{-}\varphi)$ is as in the statement of Theorem 5.1. In particular, the multiplication $m_{3}$ on $H^{\ast}$ will enjoy the symmetries described in Section 3.
We also define the tensor $R$ as
$R\colon(H^{r})^{\otimes 3}$ $\to H^{r}$
(5.1) $x\otimes y\otimes z$ $\mapsto\psi\big(m_{3}(x,z,y)\big),$
where $\psi\colon H^{3r-1}\xrightarrow{\sim}H^{r}$ is the inverse of $\varphi\cdot$-. In other words, $R$ is defined by the equation
$\varphi\cdot R(x,y,z)=m_{3}(x,z,y).$
Notice the swapping of $y$ and $z$ in the right hand side, which was motivated by our comparing the symmetry of the Bianchi-Massey tensor invented by Crowley–Nordström [10] with the symmetry of the “curvature” tensor $R$ (5.2). Using the nondegenerate pairing $\langle-,-\rangle_{\varphi}$ on $H^{r}$, we can look at $R$ as a multilinear map ${H^{r}}^{\otimes 4}\to{\mathbb{R}}$. With this in mind, we write $R_{ijkl}$ for the structure constants of $R$ with respect to a linear basis $(e_{i})$ of $H^{r}$:
$R_{ijkl}=\langle R(e_{i},e_{j},e_{k}),e_{l}\rangle_{\varphi}=\int\varphi\cdot R(e_{i},e_{j},e_{k})\cdot e_{l}=\int m_{3}(e_{i},e_{k},e_{j})\cdot e_{l},$
i.e.,
(5.2) $R_{ijkl}=\langle m_{3}(e_{i},e_{k},e_{j}),e_{l}\rangle.$
Lemma 5.2.
We have
(5.3) $R_{ijkl}=-(-1)^{r}R_{klji}.$
Proof.
By (3.3) we have
$R_{ijkl}=\langle m_{3}(e_{i},e_{k},e_{j}),e_{l}\rangle=-(-1)^{r}\langle m_{3}(e_{k},e_{j},e_{l}),e_{i}\rangle=-(-1)^{r}R_{klji}.$
Lemma 5.3 (Symmetries of $R$).
a) Assume $r$ is even. Then the tensor $R$ satisfies:
  1. (1)
    $R_{ijkl}=-R_{jikl}$ (antisymmetry in the first two indices)
  2. (2)
    $R_{ijkl}=-R_{ijlk}$ (antisymmetry in the last two indices)
  3. (3)
    $R_{ijkl}=R_{klij}$ (pair-swapping symmetry)
Therefore, $R$ defines a linear map $S^{2}\!\wedge^{2}H^{r}\to{\mathbb{R}}$. If $\dim H^{r}\leq 3$, then $R$ also satisfies the Bianchi identity and so it is is an algebraic curvature tensor.
b) Assume $r$ is odd. Then the tensor $R$ satisfies:
  1. (1)
    $R_{ijkl}=R_{jikl}$ (Symmetry in first two indices)
  2. (2)
    $R_{ijkl}=R_{ijlk}$ (Symmetry in last two indices)
  3. (3)
    $R_{ijkl}=R_{klij}$ (Pair swapping symmetry)
Therefore, $\hat{F}$ defines a linear map $S^{2}S^{2}H^{r}\to{\mathbb{R}}$.
Proof.
a) Assume that $r$ is even.
To prove (1), we need to show that
$\int m_{3}(x,z,x)\cdot w\equiv 0.$
This is immediate from the generalized even symmetry (3.13). To prove (2) we use Lemma 5.2 and (1) to obtain
[…]
To prove (3), we use Lemma 5.2 and (2):
[…]
Finally, it is well known, see, e.g., [1, Chapter 1, §G, 1.107], that in dimension less or equal to 3 a tensor that $R$ satisfying symmetries (1)–(3) also satisfies the Bianchi identity.
b) Assume that $r$ is odd.
(1) The symmetry in first two indices follows from (3.7).
(2) The symmetry in the last two indices follows from (5.3), taking into account the symmetry in first two indices.
(3) Using the same proof of the last assertion of Lemma 5.3, we conclude that the cyclicity and the first two symmetries yield the pair swapping symmetry.
Case a: $r$ even. We will assume $r$ is even and $\dim H^{r}\leq 3$. From the fact $r$ is even, we have that the nondegenerate pairing $\langle-,-\rangle_{\varphi}$ is symmetric, and for emphasis and a more immediate correspondence with notation and constructions in Riemannian geometry we will write $g=\langle-,-\rangle_{\varphi}$. By looking at $R$ as a linear map
$R\colon H^{r}\otimes H^{r}\to\mathrm{End}(H^{r})$
$x\otimes y\mapsto\{z\mapsto R(x,z,y)\}$
we can consider the Ricci tensor
(5.4) $\mathrm{Ric}:=\mathrm{tr}(R)\colon H^{r}\otimes H^{r}\to{\mathbb{R}}.$
If $(e_{i})$ is a linear basis of $H^{r}$, then
$\mathrm{Ric}_{ij}=g^{kl}R_{ikjl},$
where we have set $g_{kl}=g(e_{k},e_{l})=\langle e_{k},e_{l}\rangle_{\varphi}$. Since, under our assumptions, $R$ is an algebraic curvature tensor, and so it satisfies the Bianchi identity, the usual argument from Riemannian geometry tells us that $\mathrm{Ric}$ is symmetric: one has
$0=g^{kl}(R_{ijkl}+R_{jkil}+R_{kijl})=g^{kl}(R_{ijkl}+R_{jkll}-R_{ikjl})=\mathrm{Ric}_{ji}-\mathrm{Ric}_{ij},$
where one uses that $g^{kl}R_{ijkl}=0$ since $g^{kl}$ is symmetric in $k,l$ while $R_{ijkl}$ is antisymmetric. Moreover, since in dimension 2 and 3 the Weyl tensor vanishes, in these dimensions an algebraic curvature tensor is completely determined by its Ricci tensor. Therefore the multiplication $m_{3}$ is completely determined by $\mathrm{Ric}$. Notice that $m_{3}$ is a degree 1-1 Harrison 3-cocycle on $H^{\ast}$ and that $\mathrm{Ric}\cdot\varphi$ is a component in a degree 1-1 Harrison 2-cochain. Remarkably, as we are going to see in the next subsection, the way $\mathrm{Ric}$ determines $m_{3}$ is precisely through the Harrison boundary of (an extension of) $\mathrm{Ric}\cdot\varphi$. This in particular will show that the Harrison cohomology class of $m_{3}$ vanishes and so $\mathcal{A}$$^{\ast}$ is formal.
Case a.2. Now we shall find a Harrison coboundary for $m_{3}$ in the case $r$ is even and $b^{r}=2$.
$\phi\in C^{2,-1}(H^{*},H^{*})$ be Harrison 2-cochain. Assume $\varphi(x,y)=0$ unless $(\deg x,\deg y)=(0,0)\mod r$. By (4.3), we have
$\delta\phi(x,y,z)=x\phi(y,z)-\phi(xy,z)+\phi(x,yz)-\phi(x,y)z.$
Recalling that $m_{3}(x,y,z)=0$ unless $\deg(x,y,z)=(r,r,r)$, we see that the equation
$\delta\phi=m_{3}$
is equivalent to
$x\phi(y,z)-\phi(xy,z)+\phi(x,yz)-\phi(x,y)z=\begin{cases}m_{3}(x,y,z)&\text{if $\deg(x,y,z)=(r,r,r)$}\\ 0&\text{otherwise}.\end{cases}$
and so, since the pairing $\langle-,$-$\rangle$ is nondegenerate, to
(5.5) $\langle x\phi(y,z)-\phi(xy,z)+\phi(x,yz)-\phi(x,y)z,w\rangle=\begin{cases}\langle m_{3}(x,y,z),w\rangle&\text{if $\deg(x,y,z,w)=(r,r,r,r)$}\\ 0&\text{otherwise}.\end{cases}$
Recall that in dimension 2, an algebraic curvature tensor is expressed in terms of the metric and of its scalar curvature through the Kulkarni–Nomizu product as
[…]
i.e.,
(5.6) $R_{ijkl}=\frac{s}{2}(g_{ik}g_{jl}-g_{il}g_{jk}),$
where
$s=g^{ij}\mathrm{Ric}_{ij}$
is the scalar curvature of $R$. Analyzing Equation (5.5), taking into account the Lefschetz isomorphism: $\varphi\cdot-:H^{r}\to H^{3r}$, we make Ansatz that $\phi(x,y)=\lambda(x,y)\varphi$ for $x,y\in H^{r}$. Taking into account degree restrictions by Lemma 4.18, we make further Ansatz that $\phi(a,b)=0$ if $\deg(a,b)\not\in\{(r,r),(r,2r),(2r,r)\}$. Using (5.6), we are led to the following.
Proposition 5.4.
Let $\phi\in C^{2,-1}(H^{*},H^{*})$ be the Harrison cochain defined by the following formulas:
$\phi(x,y)$ $\coloneqq-\frac{s}{4}g(x,y)\cdot\varphi\qquad\text{if $x,y\in H^{r}$}$
$\phi(x,v)$ $\coloneqq-\frac{s}{4}\langle v,\varphi\rangle x\cdot\varphi\qquad\text{if $x\in H^{r},\,v\in H^{2r}$}$
$\phi(v,y)$ $\coloneqq-\frac{s}{4}\langle v,\varphi\rangle y\cdot\varphi\qquad\text{if $y\in H^{r},\,v\in H^{2r}$}$
$\phi(a,b)$ $\coloneqq 0\qquad\text{if $\deg(a,b)\not\in\{(r,r),(r,2r),(2r,r)\}$}$
Then $\delta\phi=m_{3}$.
Proof.
We consider the case when $\deg(x,y,z,w)=(r,r,r,r)$ first. By linearity, we only need consider the case $(x,y,z,w)=(e_{i},e_{k},e_{j},e_{l})$. In this case we have to prove that
(5.7) $\langle e_{i}\phi(e_{k},e_{j})-\phi(e_{i}e_{k},e_{j})+\phi(e_{i},e_{k}e_{j})-\phi(e_{i},e_{k})e_{j},e_{l}\rangle=R_{ijkl},$
but this is precisely equation (5.6). To address the case when $\deg(x,y,z,w)\neq(r,r,r,r)$, we notice that by degree reasons, the left hand side of (5.5) vanishes unless $\deg x+\deg y+\deg z+\deg w=4r$. Since $\mathcal{A}$$^{\ast}$ is (r1)(r-1)-connected, if none between $x,y,z,w$ is in degree zero then they all have degree at least $r$, and so if $\deg(x,y,z,w)\neq(r,r,r,r)$ we have $\deg x+\deg y+\deg z+\deg w>4r$ and the left hand side of (5.5) vanishes. When one among $x,y,z$ has degree zero (and so by linearity we may assume it is the element $1$), an immediate direct computation shows that the left hand side of (5.5) vanishes. We are therefore left with considering the case $\deg w=0$, i.e., by linearity $w=1$, with $x,y,z$ of degree at least $r$ and $\deg x+\deg y+\deg z=4r$. Since $\phi$ vanishes unless its arguments have degrees in $\{(r,r),(r,2r),(2r,r)\}$ this immediately restricts the possibilities for $\deg(x,y,z)$ in order to have a possibly nonzero left hand side to $(2r,r,r)$, $(r,2r,r)$ and $(r,r,2r)$. In the first case our equation reduces to
$\int\left(-\frac{s}{4}g(y,z)\cdot x\cdot\varphi+\frac{s}{4}\langle x,\varphi\rangle y\cdot\varphi\cdot z\right)=0,$
which is trivially verified. The other two cases are analogue.
Remark 5.5 (Bianchi identity for n=4r1n=4r-1, (r1)(r-1)-connected PDGCA).
Since $m_{3}$ is a Harrison cocycle ($\delta m_{3}=0$), for symmetric inputs in the $C_{\infty}$-algebra, the “shuffle product” relation must hold. In this degree, it reduces to the cyclic sum:
(5.8) $m_{3}(x,y,z)+m_{3}(y,z,x)+m_{3}(z,x,y)=0.$
Poincaré Duality (Cyclicity): The pairing $\langle m_{3}(x,y,z),w\rangle$ is cyclic. For $r$ even (which makes the manifold dimension 4r14r-1 odd-ish in the shifting sense), the cyclic property is:
(5.9) $\langle m_{3}(x,y,z),w\rangle=\langle m_{3}(y,z,w),x\rangle.$
For $R_{ijkl}=\langle m_{3}(e_{i},e_{k},e_{j}),e_{l}\rangle$, the combination of (5.8) and the (5.9) implies the Algebraic Bianchi Identity:
$R_{ijkl}+R_{jkil}+R_{kijl}=0.$
Case a.3. Now we shall find a Harrison coboundary for $m_{3}$ in the case that $r$ is even and $b^{r}=3$. The proof for this case is similar. We use the fact that in dimension 3 an algebraic curvature tensor $R$ is expressed in terms of the metric, of its Ricci tensor, and of the scalar curvature, through the Kulkarni–Nomizu product as
[…]
i.e.,
(5.10) $R_{ijkl}=g_{ik}\mathrm{Ric}_{jl}-g_{il}\mathrm{Ric}_{jk}+g_{jl}\mathrm{Ric}_{ik}-g_{jk}\mathrm{Ric}_{il}-\frac{s}{2}\left(g_{ik}g_{jl}-g_{il}g_{jk}\right).$
Let
$\eta\colon H^{r}\to H^{r}$
be the linear endomorphism defined by
$\mathrm{Ric}(x,y)=g(\eta(x),y).$
In terms of a linear basis $(e_{i})$ of $H^{r}$, this is
$\eta_{i}^{j}=\mathrm{Ric}_{ik}g^{kj}.$
Proposition 5.6.
Let $\phi\in C^{2,-1}(H^{*},H^{*})$ be the Harrison cochain defined as follows:
$\phi(x,y)$ $=\left(\frac{s}{4}g(x,y)-\mathrm{Ric}(x,y)\right)\cdot\varphi\qquad\text{if $x,y\in H^{r}$}$
$\phi(x,v)$ $=\langle v,\varphi\rangle(\frac{s}{4}x-\eta(x))\cdot\varphi\qquad\text{if $x\in H^{r},\,v\in H^{2r}$}$
$\phi(v,y)$ $=\langle v,\varphi\rangle(\frac{s}{4}y-\eta(y))\cdot\varphi\qquad\text{if $y\in H^{r},\,v\in H^{2r}$}$
$\phi(a,b)$ $=0\qquad\text{if $\deg(a,b)\not\in\{(r,r),(r,2r),(2r,r)\}$}$
Then $\phi$ is a Harrison coboundary for $m_{3}$.
Proof.
As in the proof of Proposition 5.4, we consider the case when $\deg(x,y,z,w)=(r,r,r,r)$ first, i.e., we have to prove equation (5.7). But this is precisely equation (5.10). To address the case when $\deg(x,y,z,w)\neq(r,r,r,r)$, we argue as in the proof of Proposition 5.4 and we only have to consider the case where $w=1$ and $\deg(x,y,z)$ belongs to $\{(2r,r,r),(r,2r,r),(r,r,2r)\}$. For $\deg(x,y,z)=(2r,r,r)$, our equation reduces to
$\langle x\phi(y,z)-\phi(xy,z)+\phi(x,yz)-\phi(x,y)z,w\rangle=0$
$\int\left((\frac{s}{4}g(y,z)-\mathrm{Ric}(y,z))\cdot x\cdot\varphi-\langle x,\varphi\rangle(\frac{s}{4}y-\eta(y))\cdot\varphi\cdot z\right)=0,$
which is trivially verified. The other two cases are analogue.
Case b: $r$ odd. In this case $b_{r}$ must be even since the bilinear form $\langle-,-\rangle_{\varphi}$ defines a symplectic form on $H^{r}$. Hence $b_{r}=2$. Consequently, $\dim(H^{r}\odot H^{r})=\dim(H^{r}\wedge H^{r})=1$. Since the pairing $\langle-,-\rangle_{\varphi}$ is nondegenerate, the multiplication $H^{r}\odot H^{r}\to H^{2r}$ is nonzero. Since $\dim(H^{r}\odot H^{r})=1$, a nonzero map stemming out of $H^{r}\odot H^{r}$ is automatically injective and so $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$. By [11, Theorem 3.10] stating that if $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$, then for the small quotient algebra $Q_{\mathrm{small}}^{*}$ of $\mathcal{A}$$^{\ast}$, we have $Q_{\mathrm{small}}^{k}=H^{k}$ also for k=2r1k=2r-1 and k=n2r+1k=n-2r+1. Hence, by [11, Theorem 3.10], we conclude that $Q_{\mathrm{small}}^{*}=H^{*}$, i.e., $\mathcal{Q}^{\ast}$ is formal. Therefore, also $\mathcal{A}$$^{\ast}$ is formal.
6. An extension of Cavalcanti formality result
In this section we shall prove an extension of Cavalcanti result on formality [5, Theorem 4]. We assume that the ground field $\mathbb{F}$} is $\mathbb{Q}$}.
Theorem 6.1.
Assume that $({\mathcal{A}}^{*},d^{-})$ is a (r1)(r-1)-connected Poincaré DGCA over a field $\mathbb{Q}$} of degree $n\leq 5r-2$ with $b^{r}\leq 2$. Let $H^{*}\coloneqq H^{*}({\mathcal{A}}^{*})$.
1) If there is a Lefschetz element $\varphi\in H^{n-2r}$ such that $\varphi\cdot-:H^{r}\to H^{n-r}$ is an isomorphism, then $\mathcal{A}$$^{*}$ is $C_{\infty}$-quasi-isomorphic to a minimal unital cyclic $C_{\infty}$-algebra whose multiplication $m_{k}$ vanishes for $k\geq 4$.
2) If, moreover, $n\leq 4r$ then $\mathcal{A}$$^{*}$ is formal.
Remark 6.2.
Theorem 6.1 (2) was stated in [5, Theorem 4] for the case of compact smooth manifolds.
Assume the condition of Theorem 6.1. Denote by $\langle-,-\rangle_{\varphi}:H^{r}\times H^{r}\to{\mathbb{F}}$ the graded symmetric bilinear form:
$\langle v,w\rangle_{\varphi}:=\int\varphi\cdot v\cdot w.$
Since $\varphi\cdot:H^{r}\to H^{n-r}$ is an isomorphism, the pairing $\langle-,-\rangle_{\varphi}$ is nondegenerate. We reformulate the first assertion of Theorem 6.1 in the following:
Proposition 6.3.
Under the assumptions of Theorem 6.1 the multiplication $m_{4}$ of the minimal unital cyclic $C_{\infty}$-algebra transfer $m$ of $({\mathcal{A}}^{*},d^{-})$ vanishes.
Proof of Proposition 6.3.
Theorem 2.12 of [13] implies that all multiplication $m_{k},k\geq 4,$ vanishes on the transferred minimal unital cyclic $C_{\infty}$-algebra transfer of $({\mathcal{A}}^{*},d^{-})$, if $n\leq 5r-3$. Thus to prove Proposition 6.3, it suffices to consider the case n=5r2n=5r-2.
Lemma 6.4.
[13, Lemma 2.16] Let $r\geq 2$ and let $\mathcal{A}$$^{\ast}$ be a (r1)(r-1)-connected Poincaré DGCA of degree $n$ admitting a Hodge homotopy and $m=(m_{3},m_{4},\ldots,)$ a unital $C_{\infty}$-algebra enhancement of $H^{*}\coloneqq H^{*}({\mathcal{A}}^{*})$. If $n\leq\ell(r-1)+3$, then the multiplication $m_{\ell-1}(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell-1})$ on $H^{*}$ is zero unless $\deg(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell-1})=(r,r,\dots,r)$. Moreover, for any $\alpha\in H^{r}$ we have $m_{\ell-1}(\alpha,\alpha,\dots,\alpha)=0$.
It follows from Lemma 6.4 that to prove Proposition 6.3 it suffices to show that for any $x,y\in H^{r}$ the following values
$m_{4}(x,x,x,y),m_{4}(y,x,x,x),m_{4}(x,y,x,x),m_{4}(x,x,y,x),$
$m_{4}(x,x,y,y),m_{4}(y,x,x,y),m_{4}(x,y,x,y)$
vanish.
As in the proof of Theorem 5.1, let us also denote by $\psi:H^{4r-2}\to H^{r}$ the inverse of $\varphi\cdot$-. We define $F:(H^{r})^{\otimes 4}\to H^{r}$ as follows. For $e_{i},e_{j},e_{k},e_{l},e_{s}\in H^{r}$ we set
(6.1) $F(e_{i},e_{j},e_{k},e_{l})\coloneqq\psi(m_{4}(e_{i},e_{j},e_{k},e_{l})).$
Equivalently,
(6.2) $m_{4}(e_{i},e_{j},e_{k},e_{l})=\varphi\cdot F(e_{i},e_{j},e_{k},e_{l}).$
Next, we set $\hat{F}:(H^{r})^{\otimes 5}\to{\mathbb{F}}$ by
$\hat{F}(i,j,k,l,s)\coloneqq\langle m_{4}(e_{i},e_{j},e_{k},e_{l}),e_{s}\rangle$
(6.3) $=\int\varphi\cdot F(e_{i},e_{j},e_{k},e_{l})\cdot e_{s}=\langle F(e_{i},e_{j},e_{k},e_{l}),e_{s}\rangle_{\varphi}.$
From the cyclicity (3.3), we have
(6.4) $\hat{F}(j,k,l,s,i)=\hat{F}(i,j,k,l,s).$
By degree reason, to prove Proposition 6.3 it suffices to verify the vanishing of $m_{4}$ on $H^{r}$, equivalently, the vanishing of $\hat{F}$ on $H^{r}$. Recall the symmetries (3.14), (3.15) derived in Lemma 3.8:
(6.5) $m_{4}(x,x,x,y)$ $=-m_{4}(y,x,x,x),$
(6.6) $m_{4}(x,y,y,z)$ $=-m_{4}(z,y,y,x).$
Step 1: Terms with four identical indices (Type 4-1). Consider terms like $\hat{F}(i,i,i,i,j)$. By Lemma 6.4, $m_{4}(x,x,x,x)=0$. Polarizing this identity (linearizing), we obtain:
$m_{4}(y,x,x,x)+m_{4}(x,y,x,x)+m_{4}(x,x,y,x)+m_{4}(x,x,x,y)=0.$
Taking the inner product with $x$:
$\hat{F}(y,x,x,x,x)+\hat{F}(x,y,x,x,x)+\hat{F}(x,x,y,x,x)+\hat{F}(x,x,x,y,x)=0.$
By cyclicity (6.4), all four terms are equal. Thus $4\hat{F}(i,i,i,i,j)=0$, implying $\hat{F}(i,i,i,i,j)=0$. By cyclicity, all permutations vanish.
Step 2: Terms with three identical indices (Type 3-2). The remaining terms involve three indices of one type and two of another (e.g., three $i$’s and two $j$’s). Up to cyclic permutation, there are only two distinct configurations:
  1. (1)
    Configuration A (aaabbaaabb): Terms like $\hat{F}(i,i,i,j,j)$. Using (6.5):
    $\langle m_{4}(i,i,i,j),j\rangle=-\langle m_{4}(j,i,i,i),j\rangle\implies\hat{F}(i,i,i,j,j)=-\hat{F}(j,i,i,i,j).$
    Using cyclicity on the RHS: $\hat{F}(j,i,i,i,j)=\hat{F}(i,i,i,j,j)$. Thus, $\hat{F}(i,i,i,j,j)=-\hat{F}(i,i,i,j,j)$, which implies $\hat{F}(i,i,i,j,j)=0$.
  2. (2)
    Configuration B (aababaabab): Terms like $\hat{F}(i,j,i,j,i)$. Consider the term $\hat{F}(j,i,i,j,i)=\langle m_{4}(j,i,i,j),i\rangle$. Using (6.6) with $x=j,y=i,z=j$:
    $m_{4}(j,i,i,j)=-m_{4}(j,i,i,j)\implies m_{4}(j,i,i,j)=0.$
    Consequently, $\hat{F}(j,i,i,j,i)=0$. By cyclicity, $\hat{F}(i,j,i,j,i)=\hat{F}(j,i,i,j,i)=0$.
Since all components vanish, $m_{4}\equiv 0$. This completes the proof of Proposition 6.3. ∎
Proof of Theorem 6.1(2).
By Theorem 5.1, it suffices to show that for $n=4r$ there exists a Harrison cochain $\phi\in C^{2,-1}(H^{*},H^{*})$ such that $m_{3}=\delta\phi$.
Let $\{e_{1},e_{2}\}$ be a basis of $H^{r}$.
Lemma 6.5.
Assume that $r$ is even. For any $x\in H^{r+1}$ and $e_{i},e_{j}\in\{e_{1},e_{2}\}$ the values $m_{3}(e_{i},x,e_{j})$ and $m_{3}(x,e_{i},e_{j})$ are defined completely by the values $m_{3}(x,e_{1},e_{1})$ and $m_{3}(x,e_{2},e_{2})$.
Proof.
From the cyclicity (3.3) one has
(6.7) $\langle m_{3}(e_{i},e_{i},e_{j}),x\rangle=-\langle m_{3}(e_{i},e_{j},x),e_{i}\rangle=\langle m_{3}(e_{j},x,e_{i}),e_{i}\rangle=-\langle m_{3}(x,e_{i},e_{i}),e_{j}\rangle.$
Using (3.11) for $r$ even, we obtain
(6.8) $\langle m_{3}(e_{i},e_{i},e_{j}),x\rangle=-\langle m_{3}(e_{j},e_{i},e_{i}),x\rangle=-\langle m_{3}(e_{i},e_{i},x),e_{j}\rangle=-\langle m_{3}(e_{i},x,e_{j}),e_{i}\rangle.$
Permuting $i$ and $j$, we obtain from (6.7) and (6.8)
(6.9) $\langle m_{3}(e_{i},x,e_{j}),e_{j}\rangle=-m_{3}(e_{j},x,e_{i}),e_{i}\rangle=\langle m_{3}(x,e_{j},e_{j}),e_{i}\rangle=\langle m_{3}(e_{i},e_{j},e_{j}),x\rangle.$
From (6.7), (6.8), (6.9) we obtain immediately Lemma 6.5 for $r$ even. ∎
Let
(6.10) $V:=\ker\langle m_{3}(e_{1},e_{1},e_{2}),-\rangle\cap\ker\langle m_{3}(e_{2},e_{2},e_{1}),-\rangle\subset H^{r+1}.$
Then $\mbox{\rm codim }V\leq 2$. Let $W\subset H^{r+1}$ be a complement of $V$ in $H^{r+1}$. The decomposition
$H^{r+1}=W\oplus V$
defines the projection $\mathrm{Proj}_{W}:H^{r+1}\to W$. It follows immediately form Lemma 6.5 and its proof the following
Proposition 6.6.
Assume the condition of Theorem 6.1(2) with $n=4r$. Let $m=(m_{3},m_{4})$ be a minimal unital cyclic $C_{\infty}$-algebra enhancement of $H^{*}$. Then the equation $\delta\phi=m_{3}$ has a solution $\phi\in C^{2,-1}(H^{*},H^{*})$.
Proof.
We seek $\phi\in C^{2,-1}_{cyc,\,unit}(H^{*},H^{*})$ such that $\delta\phi=m_{3}$. Recall that $H^{r}=\langle e_{1},e_{2}\rangle_{\mathbb{Q}}$. Since $n=4r$, we have
$m_{3}(e_{i},e_{j},e_{k})\in H^{3r-1},\qquad\phi(e_{i},e_{j})\in H^{2r-1}.$
By Lemma 6.4, $m_{3}(x,y,z)$ vanishes unless $x,y,z\in H^{r}$. Hence it suffices to solve $\delta\phi=m_{3}$ on triples in $(H^{r})^{\otimes 3}$.
Normalization. We impose the Ansatz
$\phi(a,b)=0\quad\text{whenever one of }a,b\notin H^{r}.$
Since $m_{3}$ vanishes outside $(H^{r})^{\otimes 3}$, this restriction does not affect solvability. Under this normalization, the Harrison differential simplifies: for $x,y,z\in H^{r}$,
(6.11) $\delta\phi(x,y,z)=x\phi(y,z)-\phi(x,y)z,$
because the two middle terms involve $\phi$ evaluated on elements of degree $2r$.
We now treat the two parity cases separately.
Case 1: $r$ even.
Graded commutativity implies
$m_{3}(e_{i},e_{j},e_{k})=m_{3}(e_{j},e_{i},e_{k}),\qquad\phi(e_{i},e_{j})=\phi(e_{j},e_{i}).$
Using symmetry in the first two variables and cyclicity, all values $m_{3}(e_{i},e_{j},e_{k})$ are determined by
$m_{3}(e_{1},e_{1},e_{2}),\qquad m_{3}(e_{2},e_{2},e_{1}).$
The independent unknowns for $\phi$ are
$\phi(e_{1},e_{1}),\quad\phi(e_{1},e_{2}),\quad\phi(e_{2},e_{2}),$
all in $H^{2r-1}$.
Applying (6.11), the equation $\delta\phi=m_{3}$ reduces to
(6.12) $m_{3}(e_{1},e_{1},e_{2})$ $=e_{1}\phi(e_{1},e_{2})-\phi(e_{1},e_{1})e_{2},$
(6.13) $m_{3}(e_{2},e_{2},e_{1})$ $=e_{2}\phi(e_{1},e_{2})-\phi(e_{2},e_{2})e_{1}.$
To prove that this linear system is consistent, we use the cocycle condition $\delta m_{3}=0$. For a $3$-cochain, the Harrison differential reads
$(\delta m_{3})(a,b,c,d)=a\,m_{3}(b,c,d)-m_{3}(ab,c,d)+m_{3}(a,bc,d)-m_{3}(a,b,cd)+m_{3}(a,b,c)\,d.$
If $a,b,c,d\in H^{r}$, then $ab,bc,cd\in H^{2r}$, and by Lemma 6.4 the three middle terms vanish. Thus
$(\delta m_{3})(a,b,c,d)=a\,m_{3}(b,c,d)+m_{3}(a,b,c)\,d.$
Evaluating ()(*) on $(e_{1},e_{1},e_{1},e_{2})$ gives
$0=e_{1}\,m_{3}(e_{1},e_{1},e_{2})+m_{3}(e_{1},e_{1},e_{1})\,e_{2}.$
Since $r$ is even, graded symmetry implies $m_{3}(e_{1},e_{1},e_{1})=3\,m_{3}(e_{1},e_{1},e_{2})$, and therefore
$e_{1}\,m_{3}(e_{1},e_{1},e_{2})+m_{3}(e_{1},e_{1},e_{2})\,e_{2}=0.$
Similarly, evaluating ()(*) on $(e_{2},e_{2},e_{2},e_{1})$ yields
$e_{2}\,m_{3}(e_{2},e_{2},e_{1})+m_{3}(e_{2},e_{2},e_{1})\,e_{1}=0.$
The relations (C1)–(C2) are precisely the compatibility conditions for the linear system (6.12)–(6.13). Hence the system admits a solution. Since there are three unknowns and only two independent equations, a solution $\phi$ exists. Thus $m_{3}=\delta\phi$.
Case 2: $r$ odd.
Now graded commutativity implies
$m_{3}(e_{i},e_{j},e_{k})=-\,m_{3}(e_{j},e_{i},e_{k}),\qquad\phi(e_{i},e_{j})=-\,\phi(e_{j},e_{i}).$
In particular, $m_{3}(e_{i},e_{i},e_{k})=0$, so
$m_{3}(e_{1},e_{1},e_{2})=0,\qquad m_{3}(e_{2},e_{2},e_{1})=0.$
Using cyclicity together with antisymmetry in the first two variables, one checks that all remaining values $m_{3}(e_{i},e_{j},e_{k})$ also vanish. Hence $m_{3}=0$ identically, and $\delta\phi=m_{3}$ holds with $\phi=0$.
In both parity cases, every normalized cyclic Harrison $3$-cocycle of degree 1-1 is a coboundary. This proves
$Harr^{3,-1}_{cyc,\,unit}(H^{*},H^{*})=0.$
6.1. Completion of the Proof of Theorem 6.1(2)
By Proposition 4.11, to complete the proof of Theorem 6.1(2) it suffices to show that $\tilde{\kappa}_{4}(m_{3},0,\phi_{2};0,0)]=0$ for some $\phi_{2}\in H^{2,-1}(H^{*},H^{*})$ satisfying $\delta\phi_{2}=m_{3}$.
By (4.20), $\tilde{\kappa}_{4}(m_{3},0,\phi_{2};0,0)]=[\phi_{2},m_{3}]=[\phi_{2},\delta\phi_{2}]$ = $\delta[\phi_{2},\phi_{2}]$. This completes the proof of Theorem 6.1.
7. Conclusion and final remarks
1) We have defined a complete set of invariants for the rational homotopy type of a space $X$ of finite type by introducing a stratification of isotopy modulo $k$ ($k\in\mathbb{N}^{+}$) on the set of minimal $C_{\infty}$-algebra enhancements. These are expressed via the obstruction sets $\mathcal{K}_{k}(m,m^{\prime})$, which enjoy a generalized additivity (Theorem 4.12, Proposition 4.16). While it is known that the Harrison cohomology class $m_{3}]$ corresponds to the Bianchi–Massey tensor of Crowley–Nordström, it remains an open and compelling problem to express the pentagon Massey tensor of Nagy–Nordström [33] in terms of our higher obstruction classes.
2) Our recursive approach via affine obstruction spaces provides a computable alternative to the classical moduli space description $V/G$ by Schlessinger–Stasheff [35]. While their work characterizes the moduli of homotopy types as a global quotient of an algebraic variety, we analyze the equivalence problem level-by-level through the filtered structure of the Gerstenhaber DGLA, allowing for explicit calculations at each level.
3) There is a significant conceptual overlap between our unital cyclic models and the equivariant string topology framework of Cieliebak, Hajek, and Volkov [9]. Our stratification provides the discrete algebraic foundation for the determination of $IBL_{\infty}$-structures, linking the isotopy modulo $k$ of $C_{\infty}$-algebras to the analytic propagators used in configuration space integrals.
4) Based on the vanishing of secondary obstructions in the low-dimensional cases studied, we conjecture that Theorem 6.1(1) can be strengthened, extending the dimension bound from $n\leq 5r-2$ to the critical threshold of $n\leq 6r-4$.
5) Finally, the classification of minimal $C_{\infty}$-algebra enhancements on $H^{*}(X)$ is deeply linked to the $L_{\infty}$-algebra of the automorphism group of $X$. Following the works by Berglund [2] and Buijs-Félix-Murillo [3] on the $L_{\infty}$ models of mapping spaces, we intend to explore the relationship between our Harrison obstruction classes and the rational homotopy type of $Map(X,X)$ in future work.
Acknowledgement
The author wishes to thank Università di Roma “La Sapienza” for hospitality and excellent working conditions during her visit in September 2025, where a part of this work was discussed. Shes thanks Pavel Hájek and Johannes Nordström for sending her the papers [9] and [33], respectively. She acknowledges the assistance from Google Gemini, OpenAI’s ChatGPT, and xAI’s Grok in searching several formulas in this paper. Computational assistance from xAI’s Grok in verifying the trace-free Ricci cochain, computational assistance from OpenAI’s Chatgpt in solving the Equation $\delta\phi=m_{3}$ for the proof of Proposition 6.6, and Google’s Gemini assistance in compiling Table 1 are gratefully acknowledged. Table 1 was computed with SageMath script, which was written with the help of the Google Gemini model, which we placed as Appendix at the end of this paper.
Appendix A Computational Verification of $m_{3}$
The following SageMath script was used to verify the non-vanishing of the primary obstruction $m_{3}$ for the invariant sector of the Fernández–Muñoz manifold.
Listing 1: SageMath script for $C_{\infty}$ transfer
# Setup the Rational Exterior Algebra
A.<mu, nu, theta, eta> = ExteriorAlgebra($Q$)
# Merkulov Homotopy H such that dH + Hd = Id - Pi
def H(form):
if form == mu * nu: return theta
if form == nu * mu: return -theta
return A(0)
# Computing the transferred m3 operation
def m3(a, b, c):
return H(a*b)*c + a*H(b*c)
# Basis for the invariant sector
x, y, z = mu, nu, eta
print(f"m3(x,y,x)={m3(x,y,x)}")

### v2

@@
-arXiv:2603.01219v1[math.AT] 01 Mar 2026
+Minimal Unital Cyclic $C_{\infty}$-Algebras and the Real and Rational Homotopy Type of Closed Manifolds
 Abstract.
@@
 Primary: 55P62; Secondary: 57R19, 13D03, 17B70
+Section 5 is written in collaboration with Domenico Fiorenza.
 1. Introduction
@@
 In [35] Schlessinger–Stasheff showed that the Halperin-Stasheff model leads to a representation of the set of homotopy type $(X,i:{\mathcal{H}}\cong H^{*}(X))$ as the quotient of a conical rational algebraic variety $V$ modulo a pro-unipotent group action $G$ [35].
-Here $V$ is the space of perturbations of the differential in the Tate–Jozefiak resolution of $H^{*}(X)$ and $G$ is pro-unipotent algebraic group.
-In this formulation “finding a complete set of invariants for rational homotopy types seems to be about the same order of difficulty as finding a complete set of invariants for the $G$-orbits of a variety VV… except for special cohomology rings” [35].
+Here $V$ is the space of perturbations of the differential in the Tate–Jozefiak resolution of $H^{*}(X)$ and $G$ is the pro-unipotent pro-algebraic group of automorphisms of the resolution that induce the identity on cohomology.
+(The pro-unipotency of $G$ follows from the fact that for simply connected spaces, any automorphism inducing the identity on cohomology is the exponential of a degree-0 nilpotent derivation.) In this formulation “finding a complete set of invariants for rational homotopy types seems to be about the same order of difficulty as finding a complete set of invariants for the $G$-orbits of a variety VV… except for special cohomology rings” [35].
 Recent refinements using unital cyclic $C_{\infty}$-algebras by Kajiura [25, 26] and Hamilton–Lazarev [18] have further specialized these tools for closed oriented manifolds.
@@
 Kadeishvili results [20], [22] stated that for a DGCA $\mathcal{A}$$^{*}$ there exists a minimal $C_{\infty}$-algebra enhancement $C_{\infty}{\mathcal{A}}^{*}$ of $H^{*}({\mathcal{A}}^{*})$ which is homotopy equivalent to $\mathcal{A}$$^{*}$.
-Furthermore, $C_{\infty}{\mathcal{A}}^{*}$ and $C_{\infty}{\mathcal{B}}^{*}$ are isotopic if and only if $\mathcal{A}$$^{*}$ and $\mathcal{B}$$^{*}$ are weakly equivalent.
+Furthermore, assuming that $H^{*}({\mathcal{A}}^{*})=H^{*}({\mathcal{B}}^{*})$, then $C_{\infty}{\mathcal{A}}^{*}$ and $C_{\infty}{\mathcal{B}}^{*}$ are isotopic if and only if $\mathcal{A}$$^{*}$ and $\mathcal{B}$$^{*}$ are weakly equivalent.
 We introduce a stratification of isotopy modulo $k$, for all $k\in{\mathbb{N}}^{+}$, on the set of all minimal $C_{\infty}$-algebra enhancements of a finite type graded commutative algebra (GCA) $H^{*}$ and show that two such $C_{\infty}$-algebra enhancements are isotopic if and only if they are isotopic modulo $k$ for any $k\in{\mathbb{N}}^{+}$.
@@
 In Sections 5 and 6, we give proofs of formality and almost formality theorems (Theorems 5.1, 6.1), using Hodge homotopy and the resulting symmetries of the transferred minimal unital cyclic $C_{\infty}$-algebras, proved in Section 3.
-In the last section we discuss related results and some open questions.
+In the last section 7 we discuss related results and some open questions.
+This paper also contains Appendix A where we place the SageMath scripts used to verify the Fernández-Muños example.
 The results presented in this paper concerning the case of (r1)(r-1)-connected manifolds of dimension n=4r1n=4r-1 with $b_{r}\leq 3$, as well as several foundational technical lemmas in the preliminaries, were developed in collaboration with Domenico Fiorenza.
@@
 The proofs regarding the $n=4r$ case, the higher-dimensional bounds up to 5r25r-2 and l(r1)+2l(r-1)+2, the introduction of isotopy modulo $k$ and the resulting stratification on the set of minimal $C_{\infty}$-algebra enhancements, the gauge-theoretic analysis of the higher obstruction classes are the work of the current author.
+Acknowledgement
+The author wishes to thank Università di Roma “La Sapienza” for hospitality and excellent working conditions during her visit in September 2025, where a part of this work was discussed.
+The author acknowledges the assistance of Google’s Gemini in developing the SageMath scripts used to verify the Fernández-Muñoz example in Subsection 4.5, which we place in Appendix A.
+Furthermore, the author thanks xAI’s Grok for symbolic calculations in verifying the trace-free Ricci cochain related to Proposition 5.6, and OpenAI’s ChatGPT for formula verification and symbolic calculations related to Proposition 6.6.
 2. Hodge homotopy and unital cyclic $C_{\infty}$-algebras
@@
 4.1], cf.
-[11, Definition 2.1].
+[12, Definition 2.1].
 A PDGCA $({\mathcal{A}}^{*},d)$ of degree $n$ over a field $\mathbb{F}$} is a DGCA ${\mathcal{A}}^{*}=\oplus_{k=0}^{n}{\mathcal{A}}^{k}$ whose cohomology ring $H^{*}({\mathcal{A}})$ is finite dimensional over $\mathbb{F}$} and has a fundamental class $\int\in(H^{n}({\mathcal{A}}))^{\vee}$ such that the pairing, defined by using the cup-product,
@@
 where $\cdot$] stands for the projection ${\mathcal{A}}^{n}\subseteq\ker d\to H^{n}({\mathcal{A}}^{*})$.
-The PDGCA $({\mathcal{A}}^{*},d)$ is called non-degenerate, [11, Definition 2.2], or oriented Poincaré duality algebra [28, Definition 2.1] if the pairing $\langle\cdot,\cdot\rangle$ on $\mathcal{A}$$^{*}$ is non-degenerate.
+The PDGCA $({\mathcal{A}}^{*},d)$ is called non-degenerate, [12, Definition 2.2], or oriented Poincaré duality algebra [28, Definition 2.1] if the pairing $\langle\cdot,\cdot\rangle$ on $\mathcal{A}$$^{*}$ is non-degenerate.
 A PDGCA $(H^{*},d=0)$ is called a Poincaré GCA.
@@
 Definition 2.4.
-[11, Definition 2.2], [13, Definition 2.4]. The graded subspace $\mathcal{H}^{\ast}:=\mathrm{Im}(\pi_{\mathcal{H}^{\ast}})$ is called the harmonic subspace for the Hodge homotopy $d^{-}$.
+[12, Definition 2.2], [13, Definition 2.4]. The graded subspace $\mathcal{H}^{\ast}:=\mathrm{Im}(\pi_{\mathcal{H}^{\ast}})$ is called the harmonic subspace for the Hodge homotopy $d^{-}$.
 Remark 2.5.
@@
 Examples 2.7.
-1) Our working example of a PDGCA over $\mathbb{R}$} admitting a Hodge homotopy is the de Rham complex $({\Omega}^{*}(M),d)$ of a closed smooth Riemann manifold $M$ with the harmonic subspace being the space of harmonic forms on $M$ [11, Remark 2.7 (1)], [13, Example 2.7].
-In [11, Theorem 3.10, (4.1)] Fiorenza-Kawai–Lê–Schwachhöfer proposed a simple algorithm to find a nondegenerate finite dimensional PDGCA which is weakly equivalent to $({\Omega}^{*}(M),d)$ if $b_{1}(M)=0$.
+1) Our working example of a PDGCA over $\mathbb{R}$} admitting a Hodge homotopy is the de Rham complex $({\Omega}^{*}(M),d)$ of a closed smooth Riemann manifold $M$ with the harmonic subspace being the space of harmonic forms on $M$ [12, Remark 2.7 (1)], [13, Example 2.7].
+In [12, Theorem 3.10, (4.1)] Fiorenza-Kawai–Lê–Schwachhöfer proposed a simple algorithm to find a nondegenerate finite dimensional PDGCA which is weakly equivalent to $({\Omega}^{*}(M),d)$ if $b_{1}(M)=0$.
+In the case of the de Rham complex, the pairing (2.2) coincides with the integration of the wedge product of forms, and the Hodge homotopy $d^{-}$ can be realized as $Gd^{*}$, where $G$ is the Green’s operator and $d^{*}$ is the adjoint of $d$.
 2) Our working example of a PDGCA over $\mathbb{Q}$} admitting a Hodge homotopy is the finite dimensional Poincaré duality algebra which is weakly equivalent to a given simply connected PDGCA over $\mathbb{Q}$} by [28, Theorem 1.1].
@@
 Proof.
-We prove (3.9) by induction on $k$.
-For $k=3$, equation (3.9) follows directly from (3.7).
-Assume that (3.9) holds for all indices up to (k1)(k-1).
-By (3.5) and by the the vanishing property (3.8) we have
+We prove (3.9) by induction on $k$. For $k=3$, equation (3.9) follows directly from (3.7). Assume that (3.9) holds for all indices up to (k1)(k-1). By (3.5) and by the vanishing property (3.8) we have
 $\widehat{m}_{k}(x,\ldots,x,y)$ $=(-1)^{k-1}d^{-}\widehat{m}_{k-1}(x,\ldots,x)\cdot y$
@@
 Remark 4.6 (Filtered pronilpotent DGLA).
-We note that the notion of isotopy modulo $k$, $k\in{\mathbb{N}}^{+}$, on the set of minimal $C_{\infty}$-algebra enhancements can be described using a filtration on $g_{H^{*}}$ that have been considered by Hamilton-Lazarev in [18].
-Namely we consider the following filtration
+We note that the notion of isotopy modulo $k$ ($k\in\mathbb{N}^{+}$) on the set of minimal $C_{\infty}$-algebra enhancements can be described using a filtration on the Gerstenhaber DGLA $\mathfrak{g}_{H^{*}}$ similar to those considered by Hamilton-Lazarev in [18].
+Namely, we consider the arity filtration:
 (4.13) $F^{k}\mathfrak{g}_{H^{*}}\coloneqq\prod_{n\geq k}C^{n,*}(H^{*},H^{*})$
-consists of Harrison cochains of arity $\geq k$. Then we have
-${\mathfrak{g}}_{H^{*}}=F^{1}{\mathfrak{g}}_{H^{*}}\supset F^{2}{\mathfrak{g}}_{H^{*}}\supset\ldots$
-such that
-1. $\delta(F^{p}{\mathfrak{g}}_{H^{*}})\subset F^{p}{\mathfrak{g}}_{H^{*}}$,
-2. $F^{p}{\mathfrak{g}}_{H^{*}},F^{q}{\mathfrak{g}}_{H^{*}}]\subset F^{p+q}{\mathfrak{g}}_{H^{*}}$,
-3. $\bigcap_{p}F^{p}{\mathfrak{g}}_{H^{*}}=0.$
-Furthermore, two elements $m,m^{\prime}\in{\mathfrak{g}}_{H^{*}}$ are isotopic modulo $k$ iff there exists $p\in{\mathfrak{g}}^{0}_{H^{*}}$ such that $e^{{\rm ad}_{p}}\ast(m)=m^{\prime}\mod F^{k}{\mathfrak{g}}_{H^{*}}$.
-In this filtered DGLA formalism, the relation $\sim_{k}$ is effectively a group action on the quotient ${\mathfrak{g}}_{H^{*}}/F^{k}{\mathfrak{g}}_{H^{*}}$.
-Hence, Proposition 4.5 is immediate.
-Note that Hamilton-Lazarev considered minimal $C_{\infty}$-algebra structures modulo filtration and obstructions to extendability of a minimal $C_{\infty}$-algebra structure modulo $F^{n}{\mathfrak{g}}_{H^{*}}$ to a minimal $C_{\infty}$-algebra structure modulo $F^{n+1}{\mathfrak{g}}_{H^{*}}$.
+consisting of Harrison cochains of arity $n\geq k$. This yields a descending sequence:
+$\mathfrak{g}_{H^{*}}=F^{1}\mathfrak{g}_{H^{*}}\supset F^{2}\mathfrak{g}_{H^{*}}\supset\dots$
+such that:
+  1. (1)
+    $\delta(F^{p}\mathfrak{g}_{H^{*}})\subset F^{p}\mathfrak{g}_{H^{*}}$,
+  2. (2)
+    $F^{p}\mathfrak{g}_{H^{*}},F^{q}\mathfrak{g}_{H^{*}}]\subset F^{p+q}\mathfrak{g}_{H^{*}}$,
+  3. (3)
+    $\bigcap_{p}F^{p}\mathfrak{g}_{H^{*}}=0$.
+Furthermore, two elements $m,m^{\prime}\in\mathfrak{g}_{H^{*}}$ are isotopic modulo $k$ if and only if there exists a gauge element $p\in\mathfrak{g}^{0}_{H^{*}}$ such that $e^{\text{ad}_{p}}\ast m=m^{\prime}\pmod{F^{k}\mathfrak{g}_{H^{*}}}$.
+In this filtered DGLA formalism, the relation $\sim_{k}$ is effectively the action of the pronilpotent group associated to $\mathfrak{g}_{H^{*}}$ on the quotient $\mathfrak{g}_{H^{*}}/F^{k}\mathfrak{g}_{H^{*}}$.
+Consequently, Proposition 4.5 follows immediately from the general theory of filtered DGLAs.
+We note that Hamilton-Lazarev utilized this framework to study obstructions to the extendibility of $C_{\infty}$-structures from $F^{n}$ to $F^{n+1}$.
+While the DGLA structure on Harrison cochains traces back to Gerstenhaber [14] and Voronov-Gerstenhaber [38], the emphasis on the filtered pronilpotent structure is a hallmark of the deformation-theoretic approach pioneered by Schlessinger-Stasheff [35] and Sullivan [36].
 If $H^{*}$ is a PGCA then we denote by ${\mathfrak{g}}_{H^{*},\,cycl,\,unit}$ the sub-DGLA of ${\mathfrak{g}}_{H^{*}}$ consisting of all normalized cyclic Harrison cochains.
@@
 The equivalence (2) $\iff$ (3) holds because ${\mathfrak{g}}_{H^{*},\,cycl,\,unit}$ is a filtered sub-DGLA of the filtered DGLA ${\mathfrak{g}}_{H^{*}}$. ∎
-4.3. The Primary Obstruction: Equivalence Modulo 3
+4.3. The Primary Obstruction: Isotopy Modulo 3
 The first step in comparing two $C_{\infty}$-algebra enhancements is to analyze their isotopy modulo $3$.
@@
 Assume that $m\sim_{3}m^{\prime}$.
-Then there exists an element $p=\phi_{2}+\ldots\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such $e^{{\rm ad}_{p^{\prime}}}\ast m=m\mod F^{3}{\mathfrak{g}}_{H^{*}}$.
+Then there exists an element $p=\phi_{2}+\ldots\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such $e^{{\rm ad}_{p^{\prime}}}\ast m=m\pmod{F^{3}{\mathfrak{g}}_{H^{*}}}$.
 By (4.16), we have
@@
 We are interested in the question if there exists an element $p=(\phi_{2},\phi_{3},\ldots)\in F^{2}{\mathfrak{g}}^{0}_{H^{*}}$ such that
-(4.18) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{4}{\mathfrak{g}}_{H^{*}}.$
+(4.18) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{4}{\mathfrak{g}}_{H^{*}}}$.
 In this case, we have
@@
 $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})\coloneqq$ $\{\tilde{\kappa}_{k}(p,m,m^{\prime})\coloneqq\mathrm{Proj}_{k}(e^{{\rm ad}_{p}}\ast m-m^{\prime})\in F^{k}{\mathfrak{g}}_{H^{*}}/F^{k+1}{\mathfrak{g}}_{H^{*}},$
-(4.25) $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*}},\,e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}{\mathfrak{g}}_{H^{*}}\}.$
-1) Then $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ consists of cocycles in $C^{k,(2-k)}(H^{*},H^{*})$ and the image ${\mathcal{K}}_{k}(m,m^{\prime})$ of $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ in $Harr^{k,-(k-2)}(H^{*},H^{*})$ depends only on $m\mod F^{k}=(m_{3},\ldots,m_{k})$ and $m\mod F^{k}=(m_{3}^{\prime},\ldots,m_{k}^{\prime})$.
+(4.25) $p\in F^{2}{\mathfrak{g}}^{0}_{H^{*}},\,e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{k-1}{\mathfrak{g}}_{H^{*}}}\}$.
+1) Then $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ consists of cocycles in $C^{k,(2-k)}(H^{*},H^{*})$ and the image ${\mathcal{K}}_{k}(m,m^{\prime})$ of $\widetilde{\mathcal{K}}_{k}(m,m^{\prime})$ in $Harr^{k,-(k-2)}(H^{*},H^{*})$ depends only on $m\pmod{F^{k}}=(m_{3},\ldots,m_{k})$ and $m\pmod{F^{k}}=(m_{3}^{\prime},\ldots,m_{k}^{\prime})$.
 2) Furthermore
@@
 (4.25))
-(4.28) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}{\mathfrak{g}}.$
+(4.28) $e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{k-1}{\mathfrak{g}}}$.
 Denote by
@@
 Substitute $M=m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})+R$ into the MC equation for $M$ and work modulo $F^{k+1}$:
-$\delta\big(m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big)+\frac{1}{2}\big[m^{\prime}+\tilde{\kappa}_{k}(p,m,^{\prime}),m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
+$\delta\big(m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big)+\frac{1}{2}\big[m^{\prime}+\tilde{\kappa}_{k}(p,m,^{\prime}),m^{\prime}+\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\pmod{F^{k+1}}$.
 It follows that
 $\left(\delta m^{\prime}+\frac{1}{2}[m^{\prime},m^{\prime}]\right)+\delta\tilde{\kappa}_{k}(p,m,m^{\prime})+\big[m^{\prime},\tilde{\kappa}_{k}(p,m,m^{\prime})\big]$
-$\frac{1}{2}\big[\tilde{\kappa}_{k}(p,m,m^{\prime}),\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\mod F^{k+1}.$
+$\frac{1}{2}\big[\tilde{\kappa}_{k}(p,m,m^{\prime}),\tilde{\kappa}_{k}(p,m,m^{\prime})\big]=0\pmod{F^{k+1}}$.
 The first term in the above equality vanishes since $m^{\prime}$ is a MC element.
 The last two terms vanish since $m^{\prime}\in F^{3}$ and $\tilde{\kappa}_{k}\in F^{k}$, $k\geq 3$.
-It follows that $\delta\tilde{\kappa}_{k}(p,m,m^{\prime})=0\mod F^{k+1}$.
+It follows that $\delta\tilde{\kappa}_{k}(p,m,m^{\prime})=0\pmod{F^{k+1}}$.
 This completes the proof of Lemma 4.13.
@@
 If $p,\bar{p}\in F^{2}{\mathfrak{g}}^{0}$ satisfy
-(4.31) $e^{{\rm ad}_{p}}\ast m=e^{{\rm ad}_{\bar{p}}}\ast m=m^{\prime}\mod F^{k-1}$
+(4.31) $e^{{\rm ad}_{p}}\ast m=e^{{\rm ad}_{\bar{p}}}\ast m=m^{\prime}\pmod{F^{k-1}}$
 then
@@
 Now substitute this back into the master equation (4.40)
-$(e^{\text{ad}_{q}}\ast m^{\prime})+\alpha\equiv m^{\prime}+\beta\mod{F^{k+1}}.$
+$(e^{\text{ad}_{q}}\ast m^{\prime})+\alpha\equiv m^{\prime}+\beta\pmod{{F^{k+1}}}$.
 Rearrange to isolate the gauge action on $m^{\prime}$:
-(4.41) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\equiv\beta-\alpha\mod{F^{k+1}}.$
+(4.41) $e^{\text{ad}_{q}}\ast m^{\prime}-m^{\prime}\equiv\beta-\alpha\pmod{{F^{k+1}}}$.
 Because $\alpha$ and $\beta$ are both in $F^{k}$, their difference $(\beta-\alpha)$ is also in $F^{k}$. Hence, (4.41) implies:
@@
 Lemma 4.15 (Existence of a free action).
-Let $p\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{p}}\ast m=m^{\prime}\mod F^{k-1}$. Set
+Let $p\in F^{2}{\mathfrak{g}}^{0}$ such that $e^{{\rm ad}_{p}}\ast m=m^{\prime}\pmod{F^{k-1}}$. Set
 $\tilde{\kappa}_{k}(p,m,m^{\prime})\coloneqq\mathrm{Proj}_{k}(e^{{\rm ad}_{p}}\ast m-m^{\prime})\in\tilde{\mathcal{K}}_{k}(m,m^{\prime}).$
@@
 1) Assume that $m^{\prime\prime}\in F^{3}{\mathfrak{g}}^{1}$ and $m^{\prime\prime}\sim_{k-1}m^{\prime}$. Then for any $p,q\in F^{2}{\mathfrak{g}}_{0}$ such that
-(4.51) $e^{{\rm ad}_{p}}\ast m-m^{\prime}$ $=0\mod F^{k-1},$
+(4.51) $e^{{\rm ad}_{p}}\ast m-m^{\prime}$ $=0\pmod{F^{k-1}}$,
 (4.52) $\tilde{\kappa}_{k}(p,m,m^{\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{p}}\ast m-m^{\prime}\big),$
-(4.53) $e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}$ $=0\mod F^{k-1},$
+(4.53) $e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}$ $=0\pmod{F^{k-1}}$,
 (4.54) $\tilde{\kappa}_{k}(q,m^{\prime},m^{\prime\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{q}}\ast m^{\prime}-m^{\prime\prime}\big),$
@@
 satisfies the following equations:
-(4.56) $e^{q\ast p}\ast m$ $=m^{\prime\prime}\mod F^{k-1},$
+(4.56) $e^{q\ast p}\ast m$ $=m^{\prime\prime}\pmod{F^{k-1}}$,
 (4.57) $\tilde{\kappa}_{k}(q\ast p,m,m^{\prime\prime})$ $\coloneqq\mathrm{Proj}_{k}\big(e^{{\rm ad}_{q\ast p}}\ast m-m^{\prime\prime}\big)=\tilde{\kappa}(p,m,m^{\prime})+\tilde{\kappa}(q,m^{\prime},m^{\prime\prime}).$
@@
 3) The space
-${\mathcal{K}}_{k}(p,m^{\prime},m^{\prime}):\big\{[\kappa_{k}(q,m^{\prime},m^{\prime})]\in Harr^{k,-(k-2)}(H^{*},H^{*}):$
-$q\in F^{2}{\mathfrak{g}}_{0},\,e^{{\rm ad}_{q}}\ast m^{\prime}=m\mod F^{k}\big\}$
+$\mathcal{K}_{k}(p,m^{\prime},m^{\prime})\coloneqq\left\{[\kappa_{k}(q,m^{\prime},m^{\prime})]\in\mathrm{Harr}^{k,-(k-2)}(H^{*},H^{*}):\begin{aligned} &q\in F^{2}\mathfrak{g}_{0},\\ &e^{\mathrm{ad}_{q}}\ast m^{\prime}=m\pmod{F^{k}}\end{aligned}\right\}$
 is a subgroup of the additive group $Harr^{k,-(k-2)}(H^{*},H^{*})$.
@@
 Recalling (4.53),
-$(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})=0\mod F^{k-1}.$
+$(e^{\text{ad}_{q}}*m^{\prime}-m^{\prime\prime})=0\pmod{F^{k-1}}$.
 By (4.60)
@@
 Taking into account (4.53), (4.54), we derive from (4.63) and (4.64):
-$\tilde{\kappa}_{k}(q*p,m,m^{\prime\prime})=\tilde{\kappa}_{k}(q,m,m^{\prime})+\tilde{\kappa}_{k}(p,m,m^{\prime})\mod F^{k}.$
+$\tilde{\kappa}_{k}(q*p,m,m^{\prime\prime})=\tilde{\kappa}_{k}(q,m,m^{\prime})+\tilde{\kappa}_{k}(p,m,m^{\prime})\pmod{F^{k}}$.
 This is Equation (4.57) in the first assertion of Proposition 4.16.
@@
 Taking into account Lemma 4.18, we derive Theorem 4.17 immediately from Theorems 4.7 and 4.12. ∎
-We also derive from Lemma 4.18 the following
 Corollary 4.19.
@@
 Corollary 4.19 is slightly better than [13, Theorem 2.12], which does not have “cyclic” in the formulation.
-4.5. Example: A simply connected non-formal 8-manifold
-In this subsection, we demonstrate our stratification of minimal $C_{\infty}$-algebra enhancements using the 8-dimensional simply connected symplectic manifold $\tilde{M}$ constructed by Fernández and Muñoz [12].
-This manifold serves as a crucial benchmark for our theory, as it lies on the boundary of the dimension bound $n\leq 5r-2$ for $r=2$.
-Recall that $\tilde{M}$ is obtained by a symplectic resolution of the orbifold $M/\mathbb{Z}_{3}$, where $M=\Gamma\backslash(H_{\mathbb{C}}\times\mathbb{C})$ is a compact nilmanifold.
-The Lie algebra $\mathfrak{g}$ of the complex Heisenberg group $H_{\mathbb{C}}$ has left-invariant complex 1-forms $\{\mu,\nu,\theta\}$ satisfying the structure equations:
-(4.67) $d\mu=0,\quad d\nu=0,\quad d\theta=\mu\wedge\nu.$
-The non-formality of the nilmanifold $M$ is well-known and is encoded in the non-triviality of the triple Massey product $\langle[\mu],[\nu],[\mu]\rangle$.
+4.5. Example: The Fernández-Muñoz 8-manifold
 […]
-The failure of intrinsic formality in $\tilde{M}$ can be interpreted as follows:
-  1. (1)
-    Primary Obstruction: The operation $m_{3}$ corresponds to the triple Massey product.
-In the nilmanifold $M$, the relation $d\theta=\mu\wedge\nu$ forces a non-vanishing $m_{3}([\mu],[\nu],[\cdot])$ in the transferred structure.
-  2. (2)
-    Isotopy Modulo $k$: While $\tilde{M}$ and its formal counterpart have isomorphic cohomology rings (level $k=2$), they are not isotopic modulo $3$.
-The class $m_{3}]\in Harr^{3,-1}_{cyc,\,uni}(H^{*},H^{*})=H^{4,-1}(H)$ is non-zero, hence ${\mathcal{K}}_{3}(m,0)=[m_{3}]$ where $0=(m_{3}=0,\ldots,0)$ denotes the formal enhancement.
-  3. (3)
 […]
-(In general, we have the following observation.
-Let $\tilde{M}$ be the resolution of a nilmanifold $M/G$.
-If the primary obstruction $m_{3}$ of the invariant Nomizu model $((\Lambda\mathfrak{g}^{*})^{G},d)$ is non-trivial, and the resolution map $\pi^{*}$ is a $C_{\infty}$-morphism that is injective on the generators of $m_{3}$, then $\tilde{M}$ is non-formal.)
+(4.67) $\alpha=[\mu\wedge\bar{\mu}],\quad\beta_{2}=[\nu\wedge\bar{\eta}],\quad\beta_{3}=[\bar{\nu}\wedge\eta]$
+where $\{\mu,\nu,\theta\}$ are the 1-forms of the complex Heisenberg group and $\eta$ is the fiber form.
+The structure equations $d\theta=\mu\nu$ and $d\bar{\theta}=\bar{\mu}\bar{\nu}$ imply that while these classes have trivial cup products in the cohomology ring, their higher-order interactions are non-trivial.
+Applying the Merkulov formula, the triple operation $m_{3}(\alpha,\beta_{2},\beta_{3})$ is computed as:
+$m_{3}(\alpha,\beta_{2},\beta_{3})$ $=p\left(h(i(\alpha)i(\beta_{2}))\cdot i(\beta_{3})\right)$
+$=p\left(h(-\mu\nu\bar{\mu}\bar{\eta})\cdot(\bar{\nu}\eta)\right)$
 […]
-  • $x=[\mu\wedge\bar{\eta}+\bar{\mu}\wedge\eta]$
-  • $y=[\nu\wedge\bar{\eta}+\bar{\nu}\wedge\eta]$
-  • $z=[i\eta\wedge\bar{\eta}]$
-where $\mu,\nu,\eta$ are the 1-forms of the complex Heisenberg group.
-The non-triviality of $m_{3}$ on these classes arises from the “hidden” variable $\theta$ in the relation $d\theta=\mu\wedge\nu$, which allows for the non-vanishing homotopy $H(x\wedge y)$.
-Triple $(a,b,c)$ $m_{3}(a,b,c)$ Triple $(a,b,c)$ $m_{3}(a,b,c)$
-$(x,x,y)$ $\mu\wedge\theta$ $(y,x,y)$ $2\nu\wedge\theta$
-$(x,y,x)$ $2\mu\wedge\theta$ $(y,x,z)$ $\theta\wedge\eta$
-$(x,y,y)$ $\nu\wedge\theta$ $(y,y,x)$ $\nu\wedge\theta$
-$(x,y,z)$ $\theta\wedge\eta$ $(z,x,y)$ $\theta\wedge\eta$
-$(y,x,x)$ $\mu\wedge\theta$ $(z,y,x)$ $\theta\wedge\eta$
-Table 1. Non-vanishing $m_{3}$ operations in the invariant sector of $\tilde{M}$. computed via the Merkulov formula.
+This non-vanishing $m_{3}$ represents the primary obstruction to formality.
+In our stratification, this implies that the point $\tilde{M}$ in the moduli space of homotopy types lies in the stratum where ${\mathcal{K}}_{3}(m,0)\neq 0$, distinguishing it from its formal counterpart with the same Betti numbers.
 5. A new proof of Crowley–Nordström’s formality result
@@
     $R_{ijkl}=R_{klij}$ (Pair swapping symmetry)
-Therefore, $\hat{F}$ defines a linear map $S^{2}S^{2}H^{r}\to{\mathbb{R}}$.
+Therefore, $R$ defines a linear map $S^{2}S^{2}H^{r}\to{\mathbb{R}}$.
 Proof.
@@
 Case a.2. Now we shall find a Harrison coboundary for $m_{3}$ in the case $r$ is even and $b^{r}=2$.
-$\phi\in C^{2,-1}(H^{*},H^{*})$ be Harrison 2-cochain. Assume $\varphi(x,y)=0$ unless $(\deg x,\deg y)=(0,0)\mod r$. By (4.3), we have
+$\phi\in C^{2,-1}(H^{*},H^{*})$ be Harrison 2-cochain. Assume $\varphi(x,y)=0$ unless $(\deg x,\deg y)=(0,0)\pmod{r}$. By (4.3), we have
 $\delta\phi(x,y,z)=x\phi(y,z)-\phi(xy,z)+\phi(x,yz)-\phi(x,y)z.$
@@
 Since $\dim(H^{r}\odot H^{r})=1$, a nonzero map stemming out of $H^{r}\odot H^{r}$ is automatically injective and so $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$.
-By [11, Theorem 3.10] stating that if $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$, then for the small quotient algebra $Q_{\mathrm{small}}^{*}$ of $\mathcal{A}$$^{\ast}$, we have $Q_{\mathrm{small}}^{k}=H^{k}$ also for k=2r1k=2r-1 and k=n2r+1k=n-2r+1.
-Hence, by [11, Theorem 3.10], we conclude that $Q_{\mathrm{small}}^{*}=H^{*}$, i.e., $\mathcal{Q}^{\ast}$ is formal.
+By [12, Theorem 3.10] stating that if $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$, then for the small quotient algebra $Q_{\mathrm{small}}^{*}$ of $\mathcal{A}$$^{\ast}$, we have $Q_{\mathrm{small}}^{k}=H^{k}$ also for k=2r1k=2r-1 and k=n2r+1k=n-2r+1.
+Hence, by [12, Theorem 3.10], we conclude that $Q_{\mathrm{small}}^{*}=H^{*}$, i.e., $\mathcal{Q}^{\ast}$ is formal.
 Therefore, also $\mathcal{A}$$^{\ast}$ is formal.
@@
 The remaining terms involve three indices of one type and two of another (e.g., three $i$’s and two $j$’s).
-Up to cyclic permutation, there are only two distinct configurations:
+Because $b^{r}\leq 2$, we are only dealing with two variables, up to cyclic permutation, there are only two distinct configurations:
   1. (1)
@@
 These are expressed via the obstruction sets $\mathcal{K}_{k}(m,m^{\prime})$, which enjoy a generalized additivity (Theorem 4.12, Proposition 4.16).
-While it is known that the Harrison cohomology class $m_{3}]$ corresponds to the Bianchi–Massey tensor of Crowley–Nordström, it remains an open and compelling problem to express the pentagon Massey tensor of Nagy–Nordström [33] in terms of our higher obstruction classes.
+While it is known that the Harrison cohomology class $m_{3}]$ corresponds to the Bianchi–Massey tensor of Crowley–Nordström [13, Theorem 4.4], it remains an open and compelling problem to express the pentagon Massey tensor of Nagy–Nordström [33] in terms of our higher obstruction classes.
 2) Our recursive approach via affine obstruction spaces provides a computable alternative to the classical moduli space description $V/G$ by Schlessinger–Stasheff [35].
@@
 Following the works by Berglund [2] and Buijs-Félix-Murillo [3] on the $L_{\infty}$ models of mapping spaces, we intend to explore the relationship between our Harrison obstruction classes and the rational homotopy type of $Map(X,X)$ in future work.
-Acknowledgement
-The author wishes to thank Università di Roma “La Sapienza” for hospitality and excellent working conditions during her visit in September 2025, where a part of this work was discussed.
-Shes thanks Pavel Hájek and Johannes Nordström for sending her the papers [9] and [33], respectively.
-She acknowledges the assistance from Google Gemini, OpenAI’s ChatGPT, and xAI’s Grok in searching several formulas in this paper.
-Computational assistance from xAI’s Grok in verifying the trace-free Ricci cochain, computational assistance from OpenAI’s Chatgpt in solving the Equation $\delta\phi=m_{3}$ for the proof of Proposition 6.6, and Google’s Gemini assistance in compiling Table 1 are gratefully acknowledged.
-Table 1 was computed with SageMath script, which was written with the help of the Google Gemini model, which we placed as Appendix at the end of this paper.
 Appendix A Computational Verification of $m_{3}$
 The following SageMath script was used to verify the non-vanishing of the primary obstruction $m_{3}$ for the invariant sector of the Fernández–Muñoz manifold.
-Listing 1: SageMath script for $C_{\infty}$ transfer
-# Setup the Rational Exterior Algebra
-A.<mu, nu, theta, eta> = ExteriorAlgebra($Q$)
-# Merkulov Homotopy H such that dH + Hd = Id - Pi
+Listing 1: SageMath script for $C_{\infty}$ transfer on the FM 8-manifold
+label=lst:sage_transfer
+# 1. Define the Graded Commutative Algebra over Rational Field
+# We define generators of degree 1
+names = [’mu’, ’mubar’, ’nu’, ’nubar’, ’eta’, ’etabar’, ’theta’, ’thetabar’]
+A.<mu, mubar, nu, nubar, eta, etabar, theta, thetabar> = GradedCommutativeAlgebra($Q$)
+# 2. Define the Differential d
+# d(theta) = mu*nu, d(thetabar) = mubar*nubar, others are closed
+d = A.differential({theta: mu*nu, thetabar: mubar*nubar})
+# 3. Define the Propagator (Homotopy Operator) h
+# We define h such that d*h + h*d = id - pi.
+# Here we define the mapping of coboundaries to their primitives.
 def h(form):
-if form == mu * nu: return theta
-if form == nu * mu: return -theta
-return A(0)
-# Computing the transferred m3 operation
-def m3(a, b, c):
-return H(a*b)*c + a*H(b*c)
-# Basis for the invariant sector
-x, y, z = mu, nu, eta
-print(f"m3(x,y,x)={m3(x,y,x)}")
+if form == 0: return A(0)
+res = A(0)
+# Extract monomials and coefficients
+for mon, coeff in form.dict().items():
+# Check for the Heisenberg coboundary mu*nu
+if mu*nu in A(mon):
+res += coeff * (theta * (A(mon) / (mu*nu)))
+# Check for the conjugate coboundary mubar*nubar
+elif mubar*nubar in A(mon):
+res += coeff * (thetabar * (A(mon) / (mubar*nubar)))
+return res
+# 4. Define the Cohomology Classes (Degree 2) as per FN Theorem 3.2
+alpha = mu * mubar
+beta1 = nu * nubar
+beta2 = nu * etabar
+beta3 = nubar * eta
+# 5. Merkulov Formula for m3
+# m3(a, b, c) = p( i(a)*h(i(b)*i(c)) + h(i(a)*i(b))*i(c) )
+# For the FM manifold, the projection p is the identity on non-boundaries.
+def m3_transfer(a, b, c):
+# term1: a * h(b * c)
+# term2: h(a * b) * c
+# Note: h(harmonic * harmonic) is usually 0 unless the product is a boundary.
+val = a * h(b * c) + h(a * b) * c
+return val
+# 6. Computation and Output
+triple = (alpha, beta2, beta3)
+result = m3_transfer(*triple)
+print(”—SageMathC-infinityTransfer(FNBasis)—”)
+print(f”InputTriple:alpha,beta2,beta3”)
+print(f”ResultinH^5:{result}”)
+# 7. Verification of the m3 identity
+# Check if d(result) is zero (it must be a cycle)
+print(f”Istheresultacycle?{d(result)==0}”)